### v1

# Regularity and Tailored Regularization of Deep Neural Networks, with application to parametric PDEs in uncertainty quantification

Abstract
In this paper we consider Deep Neural Networks (DNNs) with a smooth activation function as surrogates for high-dimensional functions that are somewhat smooth but costly to evaluate. We consider the standard (non-periodic) DNNs as well as propose a new model of periodic DNNs which are especially suited for a class of periodic target functions when Quasi-Monte Carlo lattice points are used as training points. We study the regularity of DNNs by obtaining explicit bounds on all mixed derivatives with respect to the input parameters. The bounds depend on the neural network parameters as well as the choice of activation function. By imposing restrictions on the network parameters to match the regularity features of the target functions, we prove that DNNs with $N$ tailor-constructed lattice training points can achieve the generalization error (or $L_{2}$ approximation error) bound ${\tt tol}+{\mathcal{O}}(N^{-r/2})$, where ${\tt tol}\in(0,1)$ is the tolerance achieved by the training error in practice, and $r=1/p^{*}$, with $p^{*}$ being the summability exponent of a sequence that characterises the decay of the input variables in the target functions, and with the implied constant independent of the dimensionality of the input data. We apply our analysis to popular models of parametric elliptic PDEs in uncertainty quantification. In our numerical experiments, we restrict the network parameters during training by adding tailored regularization terms, and we show that for an algebraic equation mimicking the parametric PDE problems the DNNs trained with tailored regularization perform significantly better.
Keywords: Deep Neural Networks, regularization, quasi-Monte Carlo methods, lattice rules, high-dimensional function approximation, parametric PDEs
MSC2020 Subject Classification: 68T07, 65D32, 65D40, 35R60
1 Introduction
We seek an efficient way to approximate Data-to-Observables maps (DtOs) for a large class of physics-based problems of the general form
[…]
where ${\boldsymbol{y}}\in Y\subset{\mathbb{R}}^{s}$ is a vector of $s$ input parameters to a parametric PDE, and $G({\boldsymbol{y}})\in Z={\mathbb{R}}^{N_{\mathrm{obs}}}$ is a vector of observables obtained from the solution to a PDE which is a somewhat smooth function of ${\boldsymbol{y}}$. The observable may be the average of the PDE solution over a subset of the domain, in which case $N_{\mathrm{obs}}=1$; or it may be the vector of values of the solution $u$ at the points of a finite element mesh, in which case $N_{\mathrm{obs}}$ may be the number of interior finite element mesh points.
Problems of this kind can be extremely challenging, especially when a high cost for one PDE solution is combined with hundreds of uncertain input parameters. Recent reviews of parametric PDEs [3, 7] both refer to the “curse of dimensionality” for large-scale computations of this kind. Thus a major theme of recent international research for parametric PDEs has been the search for fast surrogates. This research comes under many different headings, including best $N$-term approximation [8] and reduced order modelling [12, 13]. Multi-level methods [14] and multi-fidelity methods [42] play a big role, too.
This paper analyzes a popular approach to approximate $G({\boldsymbol{y}})$, namely, via a Deep Neural Network (DNN) [15, 32]. We interpret DNN mathematically as a nonlinear algorithm for multivariate function approximation, and we aim to derive conditions on the DNN to obtain favorable theoretical error bounds. Indeed, in the search for faster and more reliable methods for parametric PDEs, use is increasingly being made of machine learning [16, 38, 47], under the guiding principle that the neural network should be “physics informed” – the physics underlying the PDE system should be respected [36, 44]. We follow the papers [33, 34, 37] which used Quasi-Monte Carlo (QMC) methods to “train” the DNN, and in assuming that the training points can be chosen freely by the user, as for example in synthetic data generation (SDG). However, there are two crucial differences from [33, 34, 37]. Firstly, in addition to the standard (non-periodic) DNNs, we propose a new model of periodic DNNs and we use QMC points that are adapted to the periodicity. Secondly, we obtain explicit regularity bounds on both models of DNNs, and we impose restrictions on the neural network parameters to match the regularity features of the target functions.
1.1 The framework
We consider a standard (“feed-forward”) DNN of “depth” $L$ as an approximation of $G$ in (1) of the form
$G_{\theta}^{[L]}({\boldsymbol{y}}):=W_{L}\,\sigma(\cdots W_{2}\,\sigma(W_{1}\,\sigma(W_{0}\,{\boldsymbol{y}}+{\boldsymbol{v}}_{0})+{\boldsymbol{v}}_{1})+{\boldsymbol{v}}_{2}\cdots)+{\boldsymbol{v}}_{L},\qquad{\boldsymbol{y}}\in Y,$ (2)
where for each $\ell=0,\ldots,L$ there is an affine transformation with a “weight” matrix $W_{\ell}\in{\mathbb{R}}^{d_{\ell+1}\times d_{\ell}}$ and a “bias” vector ${\boldsymbol{v}}_{\ell}\in{\mathbb{R}}^{d_{\ell+1}\times 1}$, and there is a scalar nonlinear “activation function” $\sigma$ to be applied component-wise between the “layers”. Layer 00 is the “input” layer, with $d_{0}=s$ the dimensionality of the input vector ${\boldsymbol{y}}$. Layer $L+1$ is the “output” layer, with dimensionality $d_{L+1}=N_{\rm obs}$ being the number of observables. The $L$ layers in between are the “hidden” layers. The values of $d_{0},\ldots,d_{L+1}$ are referred to as the “widths”. This describes a general fully-connected neural network (multi-layer perceptron), with $N_{\rm DNN}:=\sum_{\ell=0}^{L}(d_{\ell+1}\times d_{\ell})+\sum_{\ell=0}^{L}d_{\ell+1}$ network parameters to be “learned” — these are collectively denoted by
$\theta:=(W_{\ell},{\boldsymbol{v}}_{\ell})_{\ell=0}^{L},$
and the set of all permissible network parameters is denoted by $\Theta$. The ensemble of depth $L$, the widths $d_{\ell}$, and other parameters used in the “training” phase (see below) are known as the “hyperparameters”. Other specific types of neural networks are defined by imposing constraints on the structure of the matrices. For example, some networks assume that all matrices on different layers are of the same size/depth, and some convolutional neural networks assume that the matrices are circulant so that the fast Fourier transform can be used to speed up the calculation. Popular activation functions include e.g., rectified linear unit (ReLU) $\sigma(x)=\max\{x,0\}$, logistic sigmoid $\sigma(x)=1/(1+e^{-x})$, hyperbolic tangent $\sigma(x)=\tanh(x)=(e^{x}-e^{-x})/(e^{x}+e^{-x})$, and swish $\sigma(x)=x/(1+e^{-x})$.
We have chosen to index $W_{\ell}$ and ${\boldsymbol{v}}_{\ell}$ starting unconventionally from $\ell=0$ for the benefit of our later presentation, with a minor pecularity that ${\boldsymbol{v}}_{\ell}$ has dimension $d_{\ell+1}$. Our definition of depth $L$ is the number of hidden layers, which is also the number of times the activation function $\sigma$ is applied. Each hidden layer $\ell$ is given by $\sigma(W_{\ell-1}\,{\boldsymbol{y}}+{\boldsymbol{v}}_{\ell-1})$ for ${\boldsymbol{y}}\in{\mathbb{R}}^{d_{\ell}\times 1}$. An interesting modification to (2), which we leave for future investigation, is to add an identity “skip link” [20, 21, 18] so that layer $\ell$ is given instead by ${\boldsymbol{y}}+\sigma(W_{\ell-1}\,{\boldsymbol{y}}+{\boldsymbol{v}}_{\ell-1})$ for ${\boldsymbol{y}}\in{\mathbb{R}}^{d_{\ell}\times 1}$. Skip links are known to be essential for training deep networks, since they avoid vanishing and exploding gradients [22].
A training set is a finite set of points ${\boldsymbol{y}}_{1},\ldots,{\boldsymbol{y}}_{N}\in Y$ which will be used together with matching (approximate) values of the observables $G({\boldsymbol{y}}_{1}),\ldots,G({\boldsymbol{y}}_{N})$ to “train” a DNN, i.e., to determine suitable network parameters $\theta$ to ensure that $G_{\theta}^{[L]}({\boldsymbol{y}})$ is a good approximation to $G({\boldsymbol{y}})$, especially in points that were not present in the training set. We refer to the pairs $\{({\boldsymbol{y}}_{k},G({\boldsymbol{y}}_{k}))\}_{k=1}^{N}$ as our training data.
We follow [33, 34, 37] in assuming that the parametric input data is appropriately scaled to the box
$Y:=[0,1]^{s}$
of Lebesgue measure $|Y|=1$, and use deterministic points ${\boldsymbol{y}}_{1},\ldots,{\boldsymbol{y}}_{N}$ as training data for a DNN instead of uniformly distributed random points. To say that the training set ${\boldsymbol{y}}_{1},\ldots,{\boldsymbol{y}}_{N}$ is a set of QMC points is to say that these points are the points of an equal-weight quadrature rule that approximates the integral over the unit cube: if $F$ is a continuous function in the unit cube then $\int_{Y}F({\boldsymbol{y}})\,{\mathrm{d}}{\boldsymbol{y}}\approx\frac{1}{N}\sum_{k=1}^{N}F({\boldsymbol{y}}_{k})$.
A DNN is “trained” by minimising an appropriate “loss function”. While our theory developed in later sections allows for an arbitrary number of observables $N_{\rm obs}$, for simplicity in the introduction we only illustrate the case $N_{\rm obs}=1$, and we consider the loss function
${\mathcal{J}}:=\frac{1}{N}\sum_{k=1}^{N}\big{(}G({\boldsymbol{y}}_{k})-G_{\theta}^{[L]}({\boldsymbol{y}}_{k})\big{)}^{2}.$
The quality of the DNN is often expressed in terms of the generalization error (or $L_{2}$ approximation error)
${\mathcal{E}}_{G}:=\bigg{(}\int_{Y}\big{(}G({\boldsymbol{y}})-G_{\theta}^{[L]}({\boldsymbol{y}})\big{)}^{2}\,{\mathrm{d}}{\boldsymbol{y}}\bigg{)}^{1/2}={\mathcal{E}}_{T}+({\mathcal{E}}_{G}-{\mathcal{E}}_{T}),$
where ${\mathcal{E}}_{T}:={\mathcal{J}}^{1/2}$ is the computable training error, and $({\mathcal{E}}_{G}-{\mathcal{E}}_{T})$ is the so-called generalization gap which is related to the quadrature error for the integrand $F({\boldsymbol{y}}):=(G({\boldsymbol{y}})-G_{\theta}^{[L]}({\boldsymbol{y}}))^{2}$ and which as a result may be reduced with a well chosen quadrature point set. It is common in “supervised” physics based learning to choose independent and identically distributed (i.i.d.) random points ${\boldsymbol{y}}_{k}$ as the training points. With QMC points instead of random points [26], the convergence rate for the generalization gap (and hence the generalization error) may be improved: the papers [33, 37] used low discrepancy sequences such as the Sobol sequence, while the paper [34] used a family of high order QMC rules.
In this paper we are particularly interested in a family of QMC rules called (rank-$1$) lattice rules [49] whose points are given by the simple formula
${\boldsymbol{y}}_{k}:=\frac{k{\boldsymbol{z}}\bmod N}{N},\quad k=1,\ldots,N,$ (3)
where ${\boldsymbol{z}}$ is an integer vector of length $s$ chosen from $\{1,2,\ldots,N-1\}^{s}$. For convenience, we sometimes use the index 00 so that ${\boldsymbol{y}}_{0}\equiv{\boldsymbol{y}}_{N}={\boldsymbol{0}}$.
1.2 Our contribution
In recent works [19, 23, 24, 25] in the context of parametric PDEs in uncertainty quantification, the authors introduced a periodic model of the random field to exploit the full power of lattice rules. Inspired by these works, in this paper we propose a periodic variant of the DNN (2) by introducing an extra sine mapping applied componentwise to the input parameters: we define
$G_{\theta}^{[L]}({\boldsymbol{y}}):=W_{L}\,\sigma(\cdots W_{2}\,\sigma(W_{1}\,\sigma(W_{0}\,\sin(2\pi{\boldsymbol{y}})+{\boldsymbol{v}}_{0})+{\boldsymbol{v}}_{1})+{\boldsymbol{v}}_{2}\cdots)+{\boldsymbol{v}}_{L},\qquad{\boldsymbol{y}}\in Y.$ (4)
This nonlinear function is then “one-periodic” with respect to the input parameters ${\boldsymbol{y}}$, i.e., $G_{\theta}^{[L]}(y_{1},\ldots,y_{j-1},0,y_{j+1},\ldots,y_{s})=G_{\theta}^{[L]}(y_{1},\ldots,y_{j-1},1,y_{j+1},\ldots,y_{s})$ for all coordinate indices $j=1,\ldots,s$. The particular form (4) is designed to handle a target function $G({\boldsymbol{y}})$ that is itself a function of $\sin(2\pi{\boldsymbol{y}})$. While it may not be suitable for all periodic integrands in general, the resulting DNN is directly suitable for the quantities of interest from [19, 23, 24, 25]. It is not intended to be a universal choice for all periodic integrands. The extra sine mapping is related to a periodic activation function, sometimes known as “sinusoidal representation networks” or “Sirens”, see e.g., [48]. Henceforth in this paper, we shall refer to (4) as “the periodic DNN”, and to (2) as “the non-periodic DNN”.
Section 2 is devoted to regularity bounds. The main result of this paper is Theorem 1, which proves explicit bounds on the mixed derivatives of both the non-periodic and periodic DNNs. To ensure that these regularity bounds not only are finite but also do not grow with increasing dimensionality $s$ or increasing network depth $L$ or widths, we need to keep three important factors under control:
  1. 1.
    The columns of the matrix $W_{0}$ need to decay in vector $\infty$-norm. These are characterized by the sequence $(\beta_{j})_{j\geq 1}$, see (7) ahead. This sequence moderates the diminishing importance of the successive variables in ${\boldsymbol{y}}$. As the nominal dimension $s$ increases, we need $\beta_{j}\to 0$ as $j\to\infty$. The faster the decay, the smaller is the “effective dimension” [6].
  2. 2.
    The matrices $W_{1},\ldots,W_{L-1}$ need to be bounded in matrix $\infty$-norm. These are characterized by the sequence $(R_{\ell})_{\ell\geq 1}$, see (8) ahead. As the network increases in depth $L$ or as we consider even higher order derivatives, we prefer $R_{\ell}\leq\rho$ for some small constant $\rho>0$ to avoid their product from becoming too large. Since the matrix $\infty$-norm is the maximum absolute row sum, if the network widths increase then necessarily the matrix elements will need to decrease in magnitude. The last matrix $W_{L}$ does not need to be restricted as much as the others. In fact, the last matrix needs to account for the scale of the learned function in a setting with bounded activation functions.
  3. 3.
    The derivatives of the activation function need to be bounded. These are characterized by the sequence $(A_{n})_{n\geq 1}$, see (9) ahead. For example, the sigmoid activation function satisfies $|\sigma^{(n)}(x)|\leq n!$ which is manageable.
Both of the explicit regularity bounds in Theorem 1 are new; both depend explicitly on the choice of activation function. In Theorem 2 we state the regularity bounds in the form that covers three commonly used activation functions, namely, sigmoid, tanh and swish. We remark that these bounds take a very familiar form, similar to those of regularity bounds of solutions to parametric PDEs in many recent papers.
Section 3 is denoted to error analysis. Recall that the generalization gap $|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|$ is related to the quadrature error of an integrand, for example, $F({\boldsymbol{y}}):=(G({\boldsymbol{y}})-G_{\theta}^{[L]}({\boldsymbol{y}}))^{2}$ when $N_{\rm obs}=1$. Since the quadrature error for a QMC rule is bounded by the so-called worst case error of the point set, multiplied by the norm of the integrand in an appropriate weighted function space, we consider the norm of the DNN and the norm of the squared difference $F$. In Theorem 3 we provide bounds of the mixed-derivative-based norms for DNNs. In Theorem 4 we prove the key result that, if we restrict the network parameters of the DNN to match its derivative bounds to those of the target function, then the norm of this squared difference $F$ is bounded in effectively the same form. Theorem 5 then shows that we can construct tailored lattice points to achieve a guaranteed error bound for the generalization gap, with the implied constant independent of the dimensionality $s$ of the input data.
Specifically, for a multiindex ${\boldsymbol{\nu}}=(\nu_{1},\ldots,\nu_{s})\neq{\boldsymbol{0}}$, where each $\nu_{j}$ specifies the number of partial derivatives to be taken respect to the $j$-th variable of the input data ${\boldsymbol{y}}$, we find that the ${\boldsymbol{\nu}}$-th order mixed derivative of the non-periodic DNN (2) in all components of the observables has an upper bound of the form
$C\,|{\boldsymbol{\nu}}|!\,{\boldsymbol{b}}^{\boldsymbol{\nu}}=C\,\Big{(}\sum_{j=1}^{s}\nu_{j}\Big{)}!\,\prod_{j=1}^{s}b_{j}^{\nu_{j}},$
for some constant $C>0$ and some sequence $(b_{j})_{j\geq 1}$ matching the corresponding derivative bound for a target function. A more complicated upper bound is obtained for the mixed derivatives of the periodic DNN (4) and a periodic target function. In the latter periodic case, we obtain the generalization error bound
${\mathcal{E}}_{G}\leq{\tt tol}+{\mathcal{O}}\big{(}N^{-r/2}\big{)},\qquad r:=\frac{1}{p^{*}},$
where ${\tt tol}\in(0,1)$ is the tolerance achieved by the training error in practice, the implied constant in the big-${\mathcal{O}}$ bound is independent of the input dimensionality $s$, and $p^{*}\in(0,1)$ is the “summability exponent” satisfying $\sum_{j\geq 1}b_{j}^{p^{*}}<\infty$. A smaller $p^{*}$ means a faster decay of the sequence and a higher order of convergence.
Section 4 provides two example classes of target functions: a parametric diffusion problem and an algebraic equation mimicking the PDE problem. In Section 5 we propose a tailored regularization in order to “encourage” the network parameters to match the desired derivative bounds as prescribed in Theorem 4. Then as an early experiment we train DNNs for an algebraic example target function. Our results indicate that the DNNs trained with tailored regularization perform significantly better than those trained with only the standard $\ell_{2}$ regularization. Section 6 contains all the technical proofs of this paper. The real-valued induction proofs for the explicit regularity bounds of DNNs are new contributions. Finally, Section 7 provides a conclusion.
1.3 Related work
There has been an explosion of works on the theory of DNNs in the last few years. In this paper we focus on the scenario where the target functions are smooth, but have very many variables and are costly to evaluate. We assume that there is a sequence $(b_{j})_{j\geq 1}$ which describes the regularity features of the target function and there is a summability exponent $p^{*}\in(0,1)$ for which $\sum_{j\geq 1}b_{j}^{p^{*}}<\infty$, as mentioned above. This is very similar to the setting of the recent review [1] which assumes that the target functions are holomorphic (see e.g., [46]): there is a sequence $(b_{j})_{j\geq 1}$ which controls the anisotropy of the target functions, and the results were stated in terms of the summability exponent $p^{*}$.
As explained in [1], much of the recent research on the approximation theory of DNNs aim to establish existence theorems: they prove the expressivity or universality of DNNs, i.e., the existence of DNNs of a certain depth and width that approximate a class of functions to a prescribed accuracy. Such proofs are typically achieved by emulation, i.e., by showing that the network parameters can be handcrafted so that the DNNs become arbitrarily close to a known approximation method such as the best $N$-term approximation. For example, for the class of holomorphic functions mentioned above, there exist DNNs of width ${\mathcal{O}}(N^{2})$ and depth ${\mathcal{O}}(\log(N))$ that achieve the near-optimal generalization errors of ${\mathcal{O}}(N^{-(1/p^{*}-1/2)})$, see e.g., [1, Theorem 7.2].
There is however still a significant gap between the approximation theory of DNNs and their practical performance (see e.g., [4, 17]). Some papers tried to narrow this gap by developing practical existence theory (see e.g., [2]) which takes into account the practical training strategies in minimizing a loss function, including the existence of a good regularization parameter, see e.g., [1, Theorem 8.1]. For further related literature we refer the readers to the review [1] and the references therein.
The review [1] is devoted to the situation where the learning methods are independent of the sequence $(b_{j})_{j\geq 1}$; it is referred to as the “unknown anisotropy setting”. In contrast, in this paper we propose to make explicit use of our knowledge of the sequence $(b_{j})_{j\geq 1}$ in our construction of lattice training points as well as in the design of our practical regularization term. This is a realistic setting in the examples of parametric PDEs in uncertainty quantification, since we know explicitly how the random fields are modeled.
“Exploding and vanishing gradients” in neural network training have been recognized early on, leading to various approaches to address the training stability issue. Among these, the most prominent ones are the introduction of gating to enable a constant error flow during training [22], identity mappings [18, 20, 21] as a simplistic and most natural way of ensuring constant error flow, and numerous normalization techniques, for example, the currently most popular layer normalization [5] to normalize the inputs to a neuron.
In this paper, we are interested in understanding and developing the basic theory of multi-layer perceptrons (DNNs). Hence, we leave the investigation of architecture modification for future work and instead turn to the design of the loss function. Equipped with the regularity bounds in Section 2 and error analysis in Section 3, we can tailor a loss function that improves multi-layer perceptrons without architectural changes.
2 Regularity bounds for DNNs
2.1 Key definitions, assumptions, and notations
To establish regularity bounds for both variants of the DNNs (2) and (4), it is convenient to work with a recursive definition:
  • (a)
    The non-periodic DNN (2) can be defined recursively by
    $\begin{cases}G_{\theta}^{[0]}({\boldsymbol{y}}):=W_{0}\,{\boldsymbol{y}}+{\boldsymbol{v}}_{0},\\ G_{\theta}^{[\ell]}({\boldsymbol{y}}):=W_{\ell}\,\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}}))+{\boldsymbol{v}}_{\ell}\quad\mbox{for}\quad\ell\geq 1.\end{cases}$ (5)
  • (b)
    The periodic DNN (4) can be defined recursively by
    $\begin{cases}G_{\theta}^{[0]}({\boldsymbol{y}}):=W_{0}\,\sin(2\pi{\boldsymbol{y}})+{\boldsymbol{v}}_{0},\\ G_{\theta}^{[\ell]}({\boldsymbol{y}}):=W_{\ell}\,\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}}))+{\boldsymbol{v}}_{\ell}\quad\mbox{for}\quad\ell\geq 1.\end{cases}$ (6)
These recursive definitions split the compositions differently as compared to how hidden layers are typically defined (layer $\ell$ is given by $\sigma(W_{\ell-1}\,{\boldsymbol{y}}+{\boldsymbol{v}}_{\ell-1})$ for ${\boldsymbol{y}}\in{\mathbb{R}}^{d_{\ell}\times 1}$), but (5) and (6) are more convenient for our induction proofs later. The view of DNNs as in (5) and (6) allows one to relate the DNN layers to integral operators as in [27, Sec. 2].
In both cases we will assume that there exist positive sequences $(\beta_{j})_{j\geq 1}$, $(R_{\ell})_{\ell\geq 1}$, $(A_{n})_{n\geq 1}$ for which the following three key assumptions hold:
  1. 1.
    The columns of the matrix $W_{0}$ have bounded vector $\infty$-norms
    $\|W_{0,:,j}\|_{\infty}:=\max_{1\leq p\leq d_{1}}|W_{0,p,j}|\;\leq\;\beta_{j}\quad\mbox{for all}\quad j=1,\ldots,s.$ (7)
  2. 2.
    The matrices $W_{\ell}$ for $\ell\geq 1$ have bounded matrix $\infty$-norms (maximum absolute row sum)
    $\|W_{\ell}\|_{\infty}:=\max_{1\leq p\leq d_{\ell+1}}\sum_{q=1}^{d_{\ell}}|W_{\ell,p,q}|\;\leq\;R_{\ell}\quad\mbox{for all}\quad\ell\geq 1.$ (8)
  3. 3.
    The activation function $\sigma:{\mathbb{R}}\to{\mathbb{R}}$ is smooth and its derivatives satisfy
    $\|\sigma^{(n)}\|_{\infty}:=\sup_{x\in{\mathbb{R}}}|\sigma^{(n)}(x)|\;\leq\;A_{n}\quad\mbox{for all}\quad n\geq 1.$ (9)
The assumption (7) and (8) are analogous to the assumptions in [34, Propositions 3.2 and 3.8]. The assumption (9) allows for a fairly generic smooth activation functions, including the logistic sigmoid, tanh, and swish (see (14) ahead).
Let ${\mathbb{N}}_{0}:=\{0,1,2,\ldots\}$ be the set of nonnegative integers. For a multiindex ${\boldsymbol{\nu}}=(\nu_{j})_{j\geq 1}\in{\mathbb{N}}_{0}^{\infty}$ we define its order to be $|{\boldsymbol{\nu}}|:=\sum_{j\geq 1}\nu_{j}$, and we consider the index set :=$\mathcal{I}:=\{{\boldsymbol{\nu}}\in{\mathbb{N}}_{0}^{\infty}:|{\boldsymbol{\nu}}|<\infty\}$. We write ${\boldsymbol{m}}\leq{\boldsymbol{\nu}}$ when $m_{j}\leq\nu_{j}$ for all $j\geq 1$, and $\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}:=\prod_{j\geq 1}\binom{\nu_{j}}{m_{j}}$ using the notation of binomial coefficients. For any sequence ${\boldsymbol{b}}=(b_{j})_{j\geq 1}$ we write ${\boldsymbol{b}}^{\boldsymbol{\nu}}:=\prod_{j\geq 1}b_{j}^{\nu_{j}}$. For any subset ${\mathfrak{u}}$ of indices we write ${\boldsymbol{\nu}}_{\mathfrak{u}}=(\nu_{j})_{j\in{\mathfrak{u}}}$, $\binom{{\boldsymbol{\nu}}_{\mathfrak{u}}}{{\boldsymbol{m}}_{\mathfrak{u}}}:=\prod_{j\in{\mathfrak{u}}}\binom{\nu_{j}}{m_{j}}$, and ${\boldsymbol{b}}_{\mathfrak{u}}^{{\boldsymbol{\nu}}_{\mathfrak{u}}}:=\prod_{j\in{\mathfrak{u}}}b_{j}^{\nu_{j}}$. We define the mixed partial derivative operator with respect to the parametric variables ${\boldsymbol{y}}$ by $\partial^{\boldsymbol{\nu}}:=\prod_{j\geq 1}(\frac{\partial}{\partial y_{j}})^{\nu_{j}}$ and $\partial^{{\boldsymbol{\nu}}_{\mathfrak{u}}}:=\prod_{j\in{\mathfrak{u}}}(\frac{\partial}{\partial y_{j}})^{\nu_{j}}$.
It turns out that the regularity of the periodic DNN (6) depends on the Stirling number of the second kind. They are defined for integers $n$ and $k$ by ${\mathcal{S}}(n,0):=\delta_{n,0}$ (i.e., it is 1 if $n=0$ and is 00 otherwise), ${\mathcal{S}}(n,k):=0$ for $k>n$, and otherwise
${\mathcal{S}}(n,k):=\frac{1}{k!}\sum_{i=0}^{k}(-1)^{k-i}\binom{k}{i}i^{n},\qquad n\geq k.$
In combinatorics, ${\mathcal{S}}(n,k)$ is the number of ways to partition a set of $n$ objects into $k$ nonempty subsets. In particular, we have ${\mathcal{S}}(0,0)=1$, ${\mathcal{S}}(n,0)=0\;\forall n\geq 1$, ${\mathcal{S}}(n,1)=1\;\forall n\geq 1$, and ${\mathcal{S}}(n,n)=1\;\forall n\geq 0$. We introduce the shorthand notation
${\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}):=\prod_{j\geq 1}{\mathcal{S}}(\nu_{j},m_{j})\quad\mbox{and}\quad{\mathcal{S}}({\boldsymbol{\nu}}_{\mathfrak{u}},{\boldsymbol{m}}_{\mathfrak{u}}):=\prod_{j\in{\mathfrak{u}}}{\mathcal{S}}(\nu_{j},m_{j}).$
2.2 Regularity results
We are ready to present our regularity results. In the following, we will use square brackets $[\ell]$ in superscript to indicate all depth-dependent quantities. Our results are obtained by real-valued induction argument. As far as we are aware, the explicit regularity bounds are new.
Theorem 1 (Regularity bounds for DNNs)
Let the sequences $(\beta_{j})_{j\geq 1}$, $(R_{\ell})_{\ell\geq 1}$, $(A_{n})_{n\geq 1}$ be defined as in (7), (8), (9), respectively. For any depth $\ell\geq 1$, any component $1\leq p\leq d_{\ell+1}$, and any multiindex ${\boldsymbol{\nu}}\in{\mathcal{I}}$ with ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$, we have the following regularity bounds:
  • (a)
    The non-periodic DNN defined in (5) satisfies
    $|\partial^{\boldsymbol{\nu}}G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}|\leq R_{\ell}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\,\Gamma_{|{\boldsymbol{\nu}}|}^{[\ell]}.\qquad\qquad\qquad\qquad\quad\;\,$ (10)
  • (b)
    The periodic DNN defined in (6) satisfies
    $|\partial^{\boldsymbol{\nu}}G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}|\leq R_{\ell}\,(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}{\boldsymbol{\beta}}^{{\boldsymbol{m}}}\,\Gamma_{|{\boldsymbol{m}}|}^{[\ell]}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}).$ (11)
In both cases the sequence $\Gamma_{n}^{[\ell]}$ is defined recursively by
$\Gamma_{n}^{[1]}$ $:=A_{n}\quad\mbox{for $n\geq 1$},\quad$ $\Gamma_{n}^{[\ell]}$ $:=\sum_{\lambda=1}^{n}A_{\lambda}\,R_{\ell-1}^{\lambda}\,{\mathbb{B}}_{n,\lambda}^{[\ell-1]}\quad\mbox{for $\ell\geq 2$ and $n\geq 1$,}$ (12)
${\mathbb{B}}_{n,1}^{[\ell]}$ $:=\Gamma_{n}^{[\ell]}\quad\mbox{for $n\geq 1$},\quad$ ${\mathbb{B}}_{n,\lambda}^{[\ell]}$ $:=\sum_{i=\lambda-1}^{n-1}\binom{n-1}{i}\,\Gamma_{n-i}^{[\ell]}\,{\mathbb{B}}_{i,\lambda-1}^{[\ell]}\quad\mbox{for $\ell\geq 1$ and $n\geq\lambda\geq 2$}.$ (13)
Proof.
See Section 6.1 for the proof of (10) and Section 6.2 for the proof of (11). ∎
The regularity bounds (10) and (11) show structured dependence of the DNNs on the columns of the matrix $W_{0}$ (via the sequence $\beta_{j}$ in (7)) and the subsequent matrices $W_{\ell}$ (via the sequence $R_{\ell}$ in (8)) as well as on the activation function $\sigma$ (via the sequence $A_{n}$ in (9)). In particular, the sequence $\Gamma_{n}^{[\ell]}$ is determined by the combined effect of the activation function $\sigma$ (via the sequence $A_{n}$) and the interior matrices $W_{1},\ldots,W_{\ell-1}$ (via the numbers $R_{1},\ldots,R_{\ell-1}$ in (8)).
When the dimensionality of the input parameters $d_{0}=s$ is large, it may make sense to impose a condition that the sequence $\beta_{j}$ is ordered and decays sufficiently fast, i.e., the columns of the matrix $W_{0}$ should decay in vector $\infty$-norms. This is consistent with the accepted wisdom in contemporary high-dimensional QMC integration theory that the underlying function space setting should be “weighted” so that the parameters or the integration variables have decaying importance.
Considering mixed derivatives with respect to the input parameters ${\boldsymbol{y}}$ with high order $|{\boldsymbol{\nu}}|$, it makes sense to restrict the matrix elements of $W_{\ell}$ such that their matrix $\infty$-norms are bounded by a reasonably small constant. At the same time, we would also want to make sure that the derivatives of the activation function $\sigma$ do not grow too quickly in magnitude so that the growth of the sequence $\Gamma_{n}^{[\ell]}$ can be sufficiently compensated by the decay of the sequence $\beta_{j}$. This has the added advantage of avoiding exploding gradients [22] with respect to the network parameters $\theta$.
In Section 6.3 we consider the following activation functions and determine explicit expressions for $A_{n}$ in (9):
[…]
The ReLU activation function $\sigma(x)=\max\{x,0\}$ is not smooth and so condition (9) does not hold. However, a generalization of swish, namely $x/(1+e^{-cx})$ for $c\rightarrow\infty$ converges to the ReLU activation function. Like ReLU, the swish function is unbounded as $x\to+\infty$.
For the remainder of this paper, we will illustrate our results for activation functions whose derivatives are bounded in the sense of (9) by the common form
$A_{n}=\xi\,\tau^{n}\,n!\qquad\mbox{for some $\xi>0$ and $\tau>0$}.$ (15)
Clearly the sigmoid case corresponds to $\xi=\tau=1$, the tanh case corresponds to $\xi=1$ and $\tau=2$, and the swish case corresponds to $\xi=1.1$ and $\tau=1$.
Theorem 2 (Regularity bounds for DNNs with $A_{n}$ of the form (15))
Let the sequences $(\beta_{j})_{j\geq 1}$, $(R_{\ell})_{\ell\geq 1}$, $(A_{n})_{n\geq 1}$ be defined as in (7), (8), (9), respectively, with $A_{n}$ given by (15). For any depth $\ell\geq 1$, any component $1\leq p\leq d_{\ell+1}$, and any multiindex ${\boldsymbol{\nu}}\in{\mathcal{I}}$ ((((including ${\boldsymbol{\nu}}={\boldsymbol{0}}$)))), we have the following regularity bounds:
  • (a)
    The non-periodic DNN defined in (5) satisfies
    $|\partial^{\boldsymbol{\nu}}G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}|\leq C_{\ell}\,\,|{\boldsymbol{\nu}}|!\,(S_{\ell}\,{\boldsymbol{\beta}})^{{\boldsymbol{\nu}}}.\qquad\qquad\qquad\qquad\quad$ (16)
  • (b)
    The periodic DNN defined in (6) satisfies
    $|\partial^{\boldsymbol{\nu}}G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}|\leq C_{\ell}\,(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}|{\boldsymbol{m}}|!\,(S_{\ell}\,{\boldsymbol{\beta}})^{{\boldsymbol{m}}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}).$ (17)
In both cases we define
$C_{\ell}:=\max\bigg{\{}\|G^{[\ell]}_{\theta}\|_{\infty},\;\frac{P_{\ell}}{S_{\ell}}\bigg{\}},\quad S_{\ell}:=\tau\sum_{k=0}^{\ell-1}P_{k},\quad P_{0}:=1,\quad P_{k}:=\prod_{t=1}^{k}(\xi\,\tau\,R_{t})\quad\mbox{for $k\geq 1$}.$ (18)
Proof.
See Lemma 8 in Section 6.3 for a proof that if $A_{n}=\xi\,\tau^{n}\,n!$ then the sequence $\Gamma_{n}^{[\ell]}$ defined recursively in (12)–(13) is given explicitly for $\ell\geq 1$ by
$\Gamma_{n}^{[\ell]}$ $=P_{\ell-1}\,\bigg{(}\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{n-1}\,\xi\,\tau^{n}\,n!=(\xi\,\tau\,P_{\ell-1})\,\bigg{(}\tau\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{n-1}\,n!=\frac{\xi\,\tau\,P_{\ell-1}}{S_{\ell}}\,S_{\ell}^{n}\,n!\,.$
Substituting this into Theorem 1 and then using $R_{\ell}\,(\xi\,\tau\,P_{\ell-1})/S_{\ell}=P_{\ell}/S_{\ell}\leq C_{\ell}$ and $S_{\ell}^{|{\boldsymbol{m}}|}\,{\boldsymbol{\beta}}^{\boldsymbol{m}}=(S_{\ell}\,{\boldsymbol{\beta}})^{\boldsymbol{m}}$, we obtain the bounds (16) and (17) for ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$. The case ${\boldsymbol{\nu}}={\boldsymbol{0}}$ holds trivially since $|G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}|\leq\|G_{\theta}^{[\ell]}\|_{\infty}\leq C_{\ell}$. ∎
It is interesting to see that the regularity bounds (16) and (17) are of the POD (“product and order dependent”) and SPOD (“smoothness-driven product and order dependent”) forms which we have encountered in previous works on parametric PDEs in uncertainty quantification, see [31] and [10] where they first appeared, as well as, e.g., [19, 23, 24, 25].
Furthermore, we observe from the regularity bounds (16) and (17) that, as the depth $\ell$ increases, we would need to restrict the magnitude of the sequence $\beta_{j}$ and the values of $R_{k}$ for $k\leq\ell-1$ in order for the sum $S_{\ell}$ to be bounded.
3 Error analysis
3.1 Preliminaries
The goal of the training process in machine learning is to find the network parameters $\theta\in\Theta$ that minimize a loss function, in our case
${\mathcal{J}}:={\mathcal{J}}(\theta):=\frac{1}{N}\sum_{k=1}^{N}\|G({\boldsymbol{y}}_{k})-G_{\theta}^{[L]}({\boldsymbol{y}}_{k})\|_{2}^{2},$
where we took the vector $2$-norm with respect to the $N_{\rm obs}$ components in the observables. It is common to regularize the minimization problem and so we use
$\theta^{*}:=\operatorname*{\mathrm{arg\,min}}_{\theta\in\Theta}\,\big{(}{\mathcal{J}}(\theta)+\lambda\,{\mathcal{R}}(\theta)\big{)},$ (19)
with an appropriately chosen regularization term ${\mathcal{R}}(\theta)$.
The error of interest is the so-called generalization error (or population risk) which measures the error of the network on unseen data. This is defined as
${\mathcal{E}}_{G}:={\mathcal{E}}_{G}(\theta):=\|G-G_{\theta}^{[L]}\|_{L_{2}(Y)}=\bigg{(}\int_{Y}\|G({\boldsymbol{y}})-G_{\theta}^{[L]}({\boldsymbol{y}})\|_{2}^{2}\,{\mathrm{d}}{\boldsymbol{y}}\bigg{)}^{1/2}.$ (20)
There is also the computable training error (or empirical risk), given by
${\mathcal{E}}_{T}:={\mathcal{E}}_{T}(\theta):=\bigg{(}\frac{1}{N}\sum_{k=1}^{N}\|G({\boldsymbol{y}}_{k})-G_{\theta}^{[L]}({\boldsymbol{y}}_{k})\|_{2}^{2}\bigg{)}^{1/2},$ (21)
leading to the so-called generalization gap $|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|$. Hence we have
${\mathcal{E}}_{G}\leq{\mathcal{E}}_{T}+|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|.$ (22)
In the case that ${\mathcal{E}}_{T}\leq{\mathcal{E}}_{G}$ then we have
$\max({\mathcal{E}}_{T},{\mathcal{E}}_{G}-{\mathcal{E}}_{T})\leq{\mathcal{E}}_{G}={\mathcal{E}}_{T}+({\mathcal{E}}_{G}-{\mathcal{E}}_{T}),$
and it is the dominant one between ${\mathcal{E}}_{T}$ and $({\mathcal{E}}_{G}-{\mathcal{E}}_{T})$ that will capture the trend of ${\mathcal{E}}_{G}$.
The square roots in (20) and (21) make it difficult to analyze the generalization gap $|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|$ directly. However, we observe that the difference ${\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}$ is the quadrature error for the integrand $F({\boldsymbol{y}})=\|(G({\boldsymbol{y}})-G_{\theta}^{[L]}({\boldsymbol{y}})\|_{2}^{2}$. For example, when $N_{\rm obs}=1$, the integrand is simply $F({\boldsymbol{y}})=(G({\boldsymbol{y}})-G_{\theta}^{[L]}({\boldsymbol{y}}))^{2}$.
The paper [37] considered $N_{\rm obs}=1$ and $1$-norm instead of $2$-norm in the definitions of (20) and (21). Interpreting the generalization gap as a quadrature error for the integrand $F({\boldsymbol{y}})=|G({\boldsymbol{y}})-G_{\theta}^{[L]}({\boldsymbol{y}})|$, they used a smooth function to replace the absolute value function in the neighborhood of the origin, together with the Koksma-Hlawska inequality for low-discrepancy sequences, to conclude that $|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|\leq c\,(\log N)^{s}/N+2\delta$, where $\delta>0$ is their tolerance parameter used in mollifying the absolute value function.
The paper [34] considered $2$-norm as we have here together with higher order QMC rules to obtain
$|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|=\frac{|{\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}|}{{\mathcal{E}}_{G}+{\mathcal{E}}_{T}}\leq\frac{|{\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}|}{{\mathcal{E}}_{G}}\leq\frac{|{\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}|}{\inf_{\theta\in\Theta}{\mathcal{E}}_{G}(\theta)}\leq c_{\rm arch}\,N^{r},$ (23)
where the finiteness of $c_{\rm arch}$ relies on an additional assumption, see [34, formula (3.9)], which depends on the network architecture. An easy alternative bound in [34] is given by
$|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|\leq\sqrt{|{\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}|}\leq c\,N^{-r/2}$ (24)
with a constant $c$ independent of the network architecture, but at only half the convergence rate. In both cases, $r=1/p^{*}$ where $p^{*}$ is the summability exponent of a sequence $(b_{j})_{j\geq 1}$ characterising the holomorphic target functions, as mentioned in the introduction.
In this paper we will instead use lattice rules in both the non-periodic and periodic settings. Furthermore, instead of the holomorphic argument, we will use the real-valued regularity bounds from the previous section.
3.2 Our error bound
For an arbitrary number of observables $N_{\rm obs}=d_{L+1}$, the error expression of interest (see (23) and (24)) satisfies
$|{\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}|$ $=\bigg{|}\int_{Y}\big{\|}G({\boldsymbol{y}})-G_{\theta}^{[L]}({\boldsymbol{y}})\big{\|}_{2}^{2}\,{\mathrm{d}}{\boldsymbol{y}}-\frac{1}{N}\sum_{k=1}^{N}\big{\|}G({\boldsymbol{y}}_{k})-G_{\theta}^{[L]}({\boldsymbol{y}}_{k})\big{\|}_{2}^{2}\,\bigg{|}$
$=\bigg{|}\int_{Y}\sum_{p=1}^{N_{\rm obs}}\big{(}G({\boldsymbol{y}})_{p}-G_{\theta}^{[L]}({\boldsymbol{y}})_{p}\big{)}^{2}\,{\mathrm{d}}{\boldsymbol{y}}-\frac{1}{N}\sum_{k=1}^{N}\sum_{p=1}^{N_{\rm obs}}\big{(}G({\boldsymbol{y}}_{k})_{p}-G_{\theta}^{[L]}({\boldsymbol{y}}_{k})_{p}\big{)}^{2}\bigg{|}$
$\leq\sum_{p=1}^{N_{\rm obs}}\bigg{|}\int_{Y}\big{(}G({\boldsymbol{y}})_{p}-G_{\theta}^{[L]}({\boldsymbol{y}})_{p}\big{)}^{2}\,{\mathrm{d}}{\boldsymbol{y}}-\frac{1}{N}\sum_{k=1}^{N}\big{(}G({\boldsymbol{y}}_{k})_{p}-G_{\theta}^{[L]}({\boldsymbol{y}}_{k})_{p}\big{)}^{2}\bigg{|}$
$\leq\sum_{p=1}^{N_{\rm obs}}e^{\rm wor}_{N}({\boldsymbol{z}})\,\big{\|}\big{(}G(\cdot)_{p}-G_{\theta}^{[L]}(\cdot)_{p}\big{)}^{2}\big{\|}_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}},$ (25)
where for each component $p$ we have the worst case integration error $e^{\rm wor}_{N}({\boldsymbol{z}})$ for a lattice rule (with $N$ points and a generating vector ${\boldsymbol{z}}$) in a weighted function space ${\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}$, multiplied by the norm of the scalar-valued integrand $F({\boldsymbol{y}}):=(G({\boldsymbol{y}})_{p}-G_{\theta}^{[L]}({\boldsymbol{y}})_{p})^{2}$ in this function space. This space has an integer smoothness parameter $\alpha$ corresponding to the order of derivatives available in each variable, together with positive weight parameters ${\boldsymbol{\gamma}}=(\gamma_{\mathfrak{u}})_{{\mathfrak{u}}\subseteq\{1:s\}}$, $\gamma_{\emptyset}:=1$, which are used to moderate the relative importance between subsets of variables, see three settings in the next subsection. To apply the theory for lattice rules, we need to show that each integrand $(G(\cdot)_{p}-G_{\theta}^{[L]}(\cdot)_{p})^{2}$ indeed belongs to the appropriate weighted function space ${\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}$ and has a nice bound on its norm.
Note that after the minimization (19) the network parameters $\theta\approx\theta^{*}$ depend on $N$ and the training points, as well as other hyperparameters in the network architecture. In summary, putting (22), (24), (3.2) together, the generalization error for the trained DNN satisfies
${\mathcal{E}}_{G}\leq{\mathcal{E}}_{T}+|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|$ $\leq{\mathcal{E}}_{T}+\sqrt{|{\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}|}$
$\leq{\mathcal{E}}_{T}+\sqrt{\textstyle\sum_{p=1}^{N_{\rm obs}}e^{\rm wor}_{N}({\boldsymbol{z}})\,\|(G(\cdot)_{p}-G_{\theta}^{[L]}(\cdot)_{p})^{2}\|_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}}.$ (26)
Hence, assuming that the training error ${\mathcal{E}}_{T}$ is under control (i.e., small relative to the generalization gap), we have a guaranteed convergence with respect to $N$ from a good lattice rule, as long as the norms $\|(G(\cdot)_{p}-G_{\theta}^{[L]}(\cdot)_{p})^{2}\|_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}$ are bounded. As $N$ increases we expect $G_{\theta}^{[L]}$ to become closer to $G$ in some pointwise sense, and these norms in (3.2) could also become smaller with increasing $N$, but we do not know how to exploit this. This situation is different from a typical QMC integration problem where the integrand is fixed and independent of $N$.
3.3 Review of three theoretical settings for lattice rules
There are known algorithms (“component-by-component construction”) for finding good lattice generating vectors ${\boldsymbol{z}}$ that achieve nearly the optimal convergence rates for the worst case errors $e^{\rm wor}_{N}({\boldsymbol{z}})$ in (3.2) in appropriately paired function space settings. Sometimes we consider “randomly shifted” lattice rules and then the error bounds hold in a root-mean-square sense with respect to the random shift. We illustrate three possible theoretical settings in this paper, see e.g., [11, Theorem 5.8] for (a), [11, Theorem 5.12] for (b), and [24, Lemma 3.1] for (c). (Note that there are lots of other theoretical results for lattice rules and other QMC rules.)
  • (a)
    For a “randomly shifted” lattice rule in a (non-periodic, “unanchored”) weighted Sobolev space ${{\mathcal{W}}_{1,{\boldsymbol{\gamma}}}}$ of smoothness parameter $\alpha=1$, we set ${\boldsymbol{\nu}}_{\mathfrak{u}}=(1,\ldots,1)$ to define
    $\|F\|_{{\mathcal{W}}_{1,{\boldsymbol{\gamma}}}}^{2}:=\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{1}{\gamma_{\mathfrak{u}}}\int_{[0,1]^{|{\mathfrak{u}}|}}\bigg{|}\int_{[0,1]^{s-|{\mathfrak{u}}|}}\partial^{{\boldsymbol{\nu}}_{\mathfrak{u}}}F({\boldsymbol{y}})\,{\mathrm{d}}{\boldsymbol{y}}_{\{1:s\}\setminus{\mathfrak{u}}}\bigg{|}^{2}{\mathrm{d}}{\boldsymbol{y}}_{\mathfrak{u}}.$ (27)
    A good generating vector ${\boldsymbol{z}}$ can be constructed such that the root-mean-square worst case error satisfies
[…]
    where $\zeta(x):=\sum_{h=1}^{\infty}h^{-x}$ is the Riemann zeta function.
  • (b)
    For a lattice rule in a (periodic, Hilbert) weighted Korobov space ${\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}$ of integer smoothness parameter $\alpha\geq 1$, we set ${\boldsymbol{\nu}}_{\mathfrak{u}}=(\alpha,\ldots,\alpha)$ to define
    $\|F\|_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}^{2}:=\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{1}{\gamma_{\mathfrak{u}}}\int_{[0,1]^{|{\mathfrak{u}}|}}\bigg{|}\int_{[0,1]^{s-|{\mathfrak{u}}|}}\partial^{{\boldsymbol{\nu}}_{\mathfrak{u}}}F({\boldsymbol{y}})\,{\mathrm{d}}{\boldsymbol{y}}_{\{1:s\}\setminus{\mathfrak{u}}}\bigg{|}^{2}{\mathrm{d}}{\boldsymbol{y}}_{\mathfrak{u}}.$ (29)
    A good generating vector ${\boldsymbol{z}}$ can be constructed such that the worst case error satisfies
    $e^{\rm wor}_{N}({\boldsymbol{z}})\leq\bigg{(}\frac{2}{N}\sum_{\emptyset\neq{\mathfrak{u}}\subseteq\{1:s\}}\gamma_{\mathfrak{u}}^{\lambda}\,\bigg{[}\frac{2\zeta(2\alpha\lambda)}{(2\pi)^{2\alpha\lambda}}\bigg{]}^{|{\mathfrak{u}}|}\bigg{)}^{\frac{1}{2\lambda}}\,\quad\forall\;\lambda\in(\tfrac{1}{2\alpha},1].$ (30)
  • (c)
    For a lattice rule in a (periodic, non-Hilbert) weighted Korobov space ${\mathcal{K}}_{\alpha,{\boldsymbol{\gamma}}}$ of integer smoothness parameter $\alpha\geq 2$, we set ${\boldsymbol{\nu}}_{\mathfrak{u}}=(\alpha,\ldots,\alpha)$ and the norm (typically defined in terms of Fourier series) is bounded by
    $\|F\|_{{\mathcal{K}}_{\alpha,{\boldsymbol{\gamma}}}}\leq\max_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{1}{\gamma_{\mathfrak{u}}}\int_{[0,1]^{|{\mathfrak{u}}|}}\bigg{|}\int_{[0,1]^{s-|{\mathfrak{u}}|}}\partial^{{\boldsymbol{\nu}}_{\mathfrak{u}}}F({\boldsymbol{y}})\,{\mathrm{d}}{\boldsymbol{y}}_{\{1:s\}\setminus{\mathfrak{u}}}\bigg{|}\,{\mathrm{d}}{\boldsymbol{y}}_{\mathfrak{u}}.$ (31)
    A good generating vector ${\boldsymbol{z}}$ can be constructed such that the worst case error satisfies
    $e^{\rm wor}_{N}({\boldsymbol{z}})\leq\bigg{(}\frac{2}{N}\sum_{\emptyset\neq{\mathfrak{u}}\subseteq\{1:s\}}\gamma_{\mathfrak{u}}^{\lambda}\,\bigg{[}\frac{2\zeta(\alpha\lambda)}{(2\pi)^{\alpha\lambda}}\bigg{]}^{|{\mathfrak{u}}|}\bigg{)}^{\frac{1}{\lambda}}\,\quad\forall\;\lambda\in(\tfrac{1}{\alpha},1].$ (32)
All three results above were stated for $N$ being a power of a prime, and similar results hold for any composite $N$. We can also construct “embedded” lattice rules [9] to be used for a range of values of $N$ (e.g., from $2^{10}$ to $2^{20}$), at the price of an extra small constant factor in the error bound.
We have deliberately stated the above results so that the norm (27) for the non-periodic setting (a) matches the norm (29) in the periodic setting (b) with $\alpha=1$. The Sobolev norm with higher smoothness for non-periodic functions will involve all mixed derivatives up to order $\alpha$, unlike (29) which contains only the highest order derivative. The Hilbert Korobov norm is typically defined in terms of the decay of Fourier series with a real parameter $\alpha>1/2$, and only when $\alpha$ is an integer can the norm be written in the form (29) involving mixed derivatives. Sometimes the scaling of $\alpha/2$ is used instead of $\alpha$, and often the scaling of $(2\pi)^{2\alpha|{\mathfrak{u}}|}$ is absorbed into the weights $\gamma_{\mathfrak{u}}$. Similarly, the non-Hilbert Korobov norm is typically defined in terms of the decay of Fourier series with a real parameter $\alpha>1$, and only when $\alpha$ is an integer can the norm satisfy the bound (31) involving mixed derivatives. The corresponding worst case error is known as “$P_{\alpha}$” in the classical theory for lattice rules [49]. We have denoted this space differently as ${\mathcal{K}}_{\alpha,{\boldsymbol{\gamma}}}$ so that there is no confusion between the norms (29) and (31).
Setting (a) can achieve close to ${\mathcal{O}}(N^{-1})$ convergence. Both settings (b) and (c) can achieve close to ${\mathcal{O}}(N^{-\alpha})$ convergence, however, the requirement on the weights $\gamma_{\mathfrak{u}}$ to ensure that the implied constants are independent of the input dimensionality $s$ can be quite different. Later, we will see an advantage of the non-Hilbert setting.
One may notice that the bound (28) has an extra factor of $2^{\lambda}$ inside the square brackets compared to (30) with $\alpha=1$; this arose from the technique of averaging with respect to the random shift and working with an associated “shift-invariant” setting. We can also apply random shifting in the periodic setting and the resulting root-mean-square error bound would remain the same as (30). Random shifting is useful as it provides a practical error estimate.
3.4 Norm bounds for DNNs
We see from (3.2) and (3.2) that we need to bound the norms $\|(G(\cdot)_{p}-G_{\theta}^{[L]}(\cdot)_{p})^{2}\|_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}$. We begin by obtaining bounds on the norm of the DNN.
Theorem 3 (Norm bounds for DNNs)
Let the sequences $(\beta_{j})_{j\geq 1}$, $(R_{\ell})_{\ell\geq 1}$, $(A_{n})_{n\geq 1}$ be defined as in (7), (8), (9), respectively, with $A_{n}$ given by (15). For depth $L\geq 1$, let $C_{L}$ and $S_{L}$ be defined as in (18), and $d_{L+1}=N_{\rm obs}$.
  • (a)
    The norm (27) of every component $1\leq p\leq N_{\rm obs}$ of the non-periodic DNN (5) satisfies
    $\big{\|}G^{[L]}_{\theta}(\cdot)_{p}\big{\|}_{{\mathcal{W}}_{1,{\boldsymbol{\gamma}}}}^{2}$ $\leq C_{L}^{2}\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{1}{\gamma_{\mathfrak{u}}}\bigg{(}|{\mathfrak{u}}|!\prod_{j\in{\mathfrak{u}}}(S_{L}\,\beta_{j})\bigg{)}^{2}.$ (33)
  • (b)
    The norm (29) of every component $1\leq p\leq N_{\rm obs}$ of the periodic DNN (6) satisfies, with ${\boldsymbol{\alpha}}_{\mathfrak{u}}=(\alpha,\ldots,\alpha)$,
    $\big{\|}G^{[L]}_{\theta}(\cdot)_{p}\big{\|}_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}^{2}$ $\leq C_{L}^{2}\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{(2\pi)^{2\alpha|{\mathfrak{u}}|}}{\gamma_{\mathfrak{u}}}\bigg{(}\sum_{{\boldsymbol{m}}_{\mathfrak{u}}\leq{\boldsymbol{\alpha}}_{\mathfrak{u}}}|{\boldsymbol{m}}_{\mathfrak{u}}|!\,{\mathcal{S}}({\boldsymbol{\alpha}}_{\mathfrak{u}},{\boldsymbol{m}}_{\mathfrak{u}})\,\prod_{j\in{\mathfrak{u}}}(S_{L}\,\beta_{j})^{m_{j}}\bigg{)}^{2}.$ (34)
  • (c)
    The norm (31) of every component $1\leq p\leq N_{\rm obs}$ of the periodic DNN (6) satisfies, with ${\boldsymbol{\alpha}}_{\mathfrak{u}}=(\alpha,\ldots,\alpha)$,
    $\big{\|}G^{[L]}_{\theta}(\cdot)_{p}\big{\|}_{{\mathcal{K}}_{\alpha,{\boldsymbol{\gamma}}}}$ $\leq C_{L}\max_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{(2\pi)^{\alpha|{\mathfrak{u}}|}}{\gamma_{\mathfrak{u}}}\sum_{{\boldsymbol{m}}_{\mathfrak{u}}\leq{\boldsymbol{\alpha}}_{\mathfrak{u}}}|{\boldsymbol{m}}_{\mathfrak{u}}|!\,{\mathcal{S}}({\boldsymbol{\alpha}}_{\mathfrak{u}},{\boldsymbol{m}}_{\mathfrak{u}})\,\prod_{j\in{\mathfrak{u}}}(S_{L}\,\beta_{j})^{m_{j}}.$ (35)
Proof.
The three results follow by applying the bounds (16) and (17) in the definition of the norms (27), (29) and (31). ∎
For bounded activation functions (e.g., $\|\sigma\|_{\infty}=1$ for sigmoid and tanh), we have $\|G^{[L]}_{\theta}\|_{\infty}\leq R_{L}\,\|\sigma\|_{\infty}+\|{\boldsymbol{v}}_{L}\|_{\infty}$, where ${\boldsymbol{v}}_{L}$ is the final vector in (2) or (4). Also, we have
$\frac{P_{L}}{S_{L}}=\frac{(\xi\,\tau\,R_{L})\,P_{L-1}}{\tau\sum_{k=0}^{L-1}P_{k}}\leq\xi\,R_{L}.$
Thus $C_{L}$ defined in (18) is bounded by
$C_{L}\leq\max\{R_{L}\,\|\sigma\|_{\infty}+\|{\boldsymbol{v}}_{L}\|_{\infty},\;\xi\,R_{L}\}.$
For unbounded activation functions (e.g., swish) the DNN can be unbounded, and it will be necessary in practice to impose a stronger restriction on $R_{L}$.
We can restrict the network parameters of the DNNs to match the regularity features of the target functions. In turn this allows us to obtain a theoretical bound on the norms in (3.2).
Theorem 4 (Match regularity of DNNs to observables)
Let the sequences $(\beta_{j})_{j\geq 1}$, $(R_{\ell})_{\ell\geq 1}$, $(A_{n})_{n\geq 1}$ be defined as in (7), (8), (9), respectively, with $A_{n}$ given by (15). For depth $L\geq 1$, let $C_{L}$ and $S_{L}$ be defined as in (18), and $d_{L+1}=N_{\rm obs}$. Given a sequence $(b_{j})_{j\geq 1}$ and a constant $C>0$, we
  • restrict the elements of the matrices $W_{1},\ldots,W_{L-1}$ so that for some constant $\rho>0$,
    $R_{\ell}\leq\rho\quad\mbox{for all $\ell=1,\ldots,L-1$},$ (36)
    $\Big{(}\mbox{thus}\quad S_{L}\leq\tau\,L\quad\mbox{if}$ $\quad\xi\,\tau\,\rho=1,\quad\mbox{and}\quad S_{L}\leq\tau\frac{(\xi\,\tau\,\rho)^{L}-1}{\xi\,\tau\,\rho-1}\quad\mbox{otherwise}\Big{)},$ (37)
  • restrict the elements of the matrix $W_{0}$ so that
    $\beta_{j}\leq\frac{b_{j}}{S_{L}}\quad\mbox{for all $j=1,\ldots,s$},$ (38)
  • restrict the matrix $W_{L}$ and vector ${\boldsymbol{v}}_{L}$ so that
    $C_{L}\leq C.$ (39)
  • (a)
    Suppose a non-periodic target function $G({\boldsymbol{y}})$ satisfies the regularity bound: for all multiindices ${\boldsymbol{\nu}}\in{\mathcal{I}}$ and all components $1\leq p\leq N_{\rm obs}$,
    $|\partial^{\boldsymbol{\nu}}G({\boldsymbol{y}})_{p}|\leq C\,|{\boldsymbol{\nu}}|!\,{\boldsymbol{b}}^{{\boldsymbol{\nu}}}.$ (40)
    Then the non-periodic DNN (2) with restrictions (36)–(39) satisfies the same regularity bound (40), and the norm (27) of both functions satisfy for all components $1\leq p\leq N_{\rm obs}$,
    $\max\bigg{\{}\begin{array}[]{r}\big{\|}G(\cdot)_{p}\big{\|}_{{\mathcal{W}}_{1,{\boldsymbol{\gamma}}}}^{2},\\ \big{\|}G^{[L]}_{\theta}(\cdot)_{p}\big{\|}_{{\mathcal{W}}_{1,{\boldsymbol{\gamma}}}}^{2}\hphantom{,}\end{array}\bigg{\}}$ $\leq C^{2}\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{1}{\gamma_{\mathfrak{u}}}\bigg{(}|{\mathfrak{u}}|!\,\prod_{j\in{\mathfrak{u}}}b_{j}\bigg{)}^{2},$ (43)
    $\big{\|}\big{(}G(\cdot)_{p}-G^{[L]}_{\theta}(\cdot)_{p}\big{)}^{2}\big{\|}_{{\mathcal{W}}_{1,{\boldsymbol{\gamma}}}}^{2}$ $\leq 16\,C^{4}\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\frac{1}{\gamma_{\mathfrak{u}}}\bigg{(}(|{\mathfrak{u}}|+1)!\,\prod_{j\in{\mathfrak{u}}}b_{j}\bigg{)}^{2}.$ (44)
  • (b)
    Suppose a periodic target function $G({\boldsymbol{y}})$ satisfies the regularity bound: for all mutiindices ${\boldsymbol{\nu}}\in{\mathcal{I}}$ and all components $1\leq p\leq N_{\rm obs}$,
    $|\partial^{\boldsymbol{\nu}}G({\boldsymbol{y}})_{p}|\leq C\,(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}|{\boldsymbol{m}}|!\,{\boldsymbol{b}}^{{\boldsymbol{m}}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}).$ (45)
    Then the periodic DNN (4) with restrictions (36)–(39) satisfies the same regularity bound (45), and the norm (29) of both functions satisfy for all components $1\leq p\leq N_{\rm obs}$,
    $\max\bigg{\{}\begin{array}[]{r}\big{\|}G(\cdot)_{p}\big{\|}_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}^{2},\\ \big{\|}G^{[L]}_{\theta}(\cdot)_{p}\big{\|}_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}^{2}\hphantom{,}\end{array}\bigg{\}}$ $\leq C^{2}\!\!\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\!\!\!\!\frac{(2\pi)^{2\alpha|{\mathfrak{u}}|}}{\gamma_{\mathfrak{u}}}\bigg{(}\sum_{{\boldsymbol{m}}_{\mathfrak{u}}\leq{\boldsymbol{\alpha}}_{\mathfrak{u}}}|{\boldsymbol{m}}_{\mathfrak{u}}|!\,{\mathcal{S}}({\boldsymbol{\alpha}}_{\mathfrak{u}},{\boldsymbol{m}}_{\mathfrak{u}})\,\prod_{j\in{\mathfrak{u}}}b_{j}^{m_{j}}\!\bigg{)}^{2},$ (48)
    $\big{\|}\big{(}G(\cdot)_{p}-G^{[L]}_{\theta}(\cdot)_{p}\big{)}^{2}\big{\|}_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}^{2}$ $\leq 16\,C^{4}\!\!\sum_{{\mathfrak{u}}\subseteq\{1:s\}}\!\!\!\!\frac{(2\pi)^{2\alpha|{\mathfrak{u}}|}}{\gamma_{\mathfrak{u}}}\bigg{(}\sum_{{\boldsymbol{m}}_{\mathfrak{u}}\leq{\boldsymbol{\alpha}}_{\mathfrak{u}}}(|{\boldsymbol{m}}_{\mathfrak{u}}|+1)!\,{\mathcal{S}}({\boldsymbol{\alpha}}_{\mathfrak{u}},{\boldsymbol{m}}_{\mathfrak{u}})\,\prod_{j\in{\mathfrak{u}}}b_{j}^{m_{j}}\!\bigg{)}^{2}.$ (49)
  • (c)
    Continuing from (b), the norm (31) of both functions satisfy for all components $p$,
    $\max\bigg{\{}\begin{array}[]{r}\big{\|}G(\cdot)_{p}\big{\|}_{{\mathcal{K}}_{\alpha,{\boldsymbol{\gamma}}}},\\ \big{\|}G^{[L]}_{\theta}(\cdot)_{p}\big{\|}_{{\mathcal{K}}_{\alpha,{\boldsymbol{\gamma}}}}\hphantom{,}\end{array}\bigg{\}}$ $\leq C\max_{{\mathfrak{u}}\subseteq\{1:s\}}\!\!\frac{(2\pi)^{\alpha|{\mathfrak{u}}|}}{\gamma_{\mathfrak{u}}}\!\!\sum_{{\boldsymbol{m}}_{\mathfrak{u}}\leq{\boldsymbol{\alpha}}_{\mathfrak{u}}}|{\boldsymbol{m}}_{\mathfrak{u}}|!\,{\mathcal{S}}({\boldsymbol{\alpha}}_{\mathfrak{u}},{\boldsymbol{m}}_{\mathfrak{u}})\,\prod_{j\in{\mathfrak{u}}}b_{j}^{m_{j}},$ (52)
    $\big{\|}\big{(}G(\cdot)_{p}-G^{[L]}_{\theta}(\cdot)_{p}\big{)}^{2}\big{\|}_{{\mathcal{K}}_{\alpha,{\boldsymbol{\gamma}}}}$ $\leq 4\,C^{2}\max_{{\mathfrak{u}}\subseteq\{1:s\}}\!\!\frac{(2\pi)^{\alpha|{\mathfrak{u}}|}}{\gamma_{\mathfrak{u}}}\!\!\sum_{{\boldsymbol{m}}_{\mathfrak{u}}\leq{\boldsymbol{\alpha}}_{\mathfrak{u}}}(|{\boldsymbol{m}}_{\mathfrak{u}}|+1)!\,{\mathcal{S}}({\boldsymbol{\alpha}}_{\mathfrak{u}},{\boldsymbol{m}}_{\mathfrak{u}})\,\prod_{j\in{\mathfrak{u}}}b_{j}^{m_{j}}.$ (53)
Proof.
With the restrictions (36)–(39), it is clear that the right-hand sides of (16) and (17) are bounded from above by the right-hand sides of (40) and (45), respectively. The bounds (43), (48) and (52) then follow from the definition of the norms (27), (29) and (31). The final bounds (44), (49) and (53) can be obtained from (40) and (45) using the Liebniz product rule, see Section 6.4. ∎
We now choose weights $\gamma_{\mathfrak{u}}$ and construct lattice generating vectors in all three cases to achieve convergence rates that are independent of the input dimension $s$.
Theorem 5 (Tailored lattice training points)
Under the same conditions as in Theorem 4, assume further that there is a “summability exponent” $p^{*}\in(0,1)$ such that $\sum_{j\geq 1}b_{j}^{p^{*}}<\infty$.
  • (a)
    Construct as in Subsection 3.3$(a)$ a good generating vector ${\boldsymbol{z}}$ for a randomly-shifted lattice rule in weighted Sobolev space with weights
    $\gamma_{\mathfrak{u}}:=\bigg{(}(|{\mathfrak{u}}|+1)!\,\prod_{j\in{\mathfrak{u}}}\bigg{(}\sqrt{\frac{(2\pi)^{2\lambda}}{2\zeta(2\lambda)\,2^{\lambda}}}\,b_{j}\bigg{)}\bigg{)}^{\frac{2}{1+\lambda}},\;\lambda:=\begin{cases}\frac{1}{2-2\delta},\delta\in(0,\frac{1}{2}),&\mbox{if }p^{*}\in(0,\frac{2}{3}],\\ \frac{p^{*}}{2-p^{*}}&\mbox{if }p^{*}\in(\frac{2}{3},1),\end{cases}$ (54)
    If a non-periodic DNN (2) is trained using these lattice points, and the training error reaches the threshold ${\tt tol}\in(0,1)$, then the generalization error is bounded by
    ${\mathcal{E}}_{G}\leq{\tt tol}+{\mathcal{O}}\big{(}N^{-r/2}\big{)},\qquad r:=\min\Big{(}1-\delta,\;\frac{1}{p^{*}}-\frac{1}{2}\Big{)},\quad\delta\in(0,\tfrac{1}{2}).$ (55)
  • (b)
    Construct as in Subsection 3.3$(b)$ a good generating vector ${\boldsymbol{z}}$ for a lattice rule in weighted Hilbert Korobov space with weights
    $\gamma_{\mathfrak{u}}:=(2\pi)^{2\alpha|{\mathfrak{u}}|}\sum_{{\boldsymbol{m}}_{\mathfrak{u}}\leq{\boldsymbol{\alpha}}_{\mathfrak{u}}}\!\!\bigg{(}(|{\boldsymbol{m}}_{\mathfrak{u}}|+1)!\,\prod_{j\in{\mathfrak{u}}}\frac{b_{j}^{m_{j}}{\mathcal{S}}(\alpha,m_{j})}{\sqrt{2\zeta(2\alpha\lambda)}}\bigg{)}^{\frac{2}{1+\lambda}},\,\alpha:=\left\lfloor\tfrac{1}{p^{*}}+\tfrac{1}{2}\right\rfloor,\,\lambda:=\tfrac{p^{*}}{2-p^{*}}.$ (56)
    If a periodic DNN (4) is trained using these lattice points, and the training error reaches the threshold ${\tt tol}\in(0,1)$, then the generalization error is bounded by
    ${\mathcal{E}}_{G}\leq{\tt tol}+{\mathcal{O}}\big{(}N^{-r/2}\big{)},\qquad r:=\frac{1}{p^{*}}-\frac{1}{2}.$ (57)
  • (c)
    Construct as in Subsection 3.3$(c)$ a good generating vector ${\boldsymbol{z}}$ for a lattice rule in weighted non-Hilbert Korobov space with weights
    $\gamma_{\mathfrak{u}}:=(2\pi)^{\alpha|{\mathfrak{u}}|}\sum_{{\boldsymbol{m}}_{\mathfrak{u}}\leq{\boldsymbol{\alpha}}_{\mathfrak{u}}}(|{\boldsymbol{m}}_{\mathfrak{u}}|+1)!\,\prod_{j\in{\mathfrak{u}}}\big{(}b_{j}^{m_{j}}{\mathcal{S}}(\alpha,m_{j})\big{)},\quad\alpha:=\left\lfloor\tfrac{1}{p^{*}}\right\rfloor+1.$ (58)
    If a periodic DNN (4) is trained using these lattice points, and the training error reaches the threshold ${\tt tol}\in(0,1)$, then the generalization error is bounded by
    ${\mathcal{E}}_{G}\leq{\tt tol}+{\mathcal{O}}\big{(}N^{-r/2}\big{)},\qquad r:=\frac{1}{p^{*}}.$ (59)
In all three cases the implied constant in the big-${\mathcal{O}}$ bound is independent of $s$.
In practice, if the condition (38) is not achieved after training, i.e., if $\beta_{j}\leq b_{j}/S_{L}$ is not true for some $j$, with $S_{L}$ defined in (18), but instead there is a constant $\kappa\geq 1/s_{L}$ such that
$\beta_{j}\leq\kappa\,b_{j}\quad\mbox{for all }j=1,\ldots,s,$ (60)
then the bounds (44), (49), (53) hold with each $b_{j}$ replaced by $\kappa\,S_{L}\,b_{j}$. Even though the lattice training points were constructed with the weights (54), (56), (58) based on the original $b_{j}$, the big-${\mathcal{O}}$ bounds (55), (57), (59) hold now with enlarged implied constants that depend on $\kappa\,S_{L}$. In particular, the enlarged constants for (55) and (57) are still bounded independently of $s$, but the enlarged constant for (59) grows exponentially with $s$.
Proof.
The results follow from (3.2) by combining (44) with (28), (49) with (30), and (53) with (32), and then choosing the weights and other parameters such as $\alpha$ and $\lambda$ with respect to $p^{*}$ to ensure that the implied constant is finite in each case. The derivations of weights and convergence rates follow the same arguments as in several recent papers, see e.g., [31, 30] for (a), [23] for (b), and [24] for (c).
The final statement regarding dimension independence for the enlarged constants in the settings (a) and (b) is true because the summability exponent is unchanged, i.e., $\sum_{j\geq 1}(\kappa\,S_{L}\,b_{j})^{p^{*}}=(\kappa\,S_{L})^{p^{*}}\sum_{j\geq 1}b_{j}^{p^{*}}<\infty$. The setting (c) is different and will incur an extra factor $(\kappa\,S_{L})^{\alpha\,s}$. Detailed derivations for all three settings are given in our survey article [28]. ∎
The theoretical convergence rates above are very conservative because we were aiming for implied constants that do not depend on $s$. Moreover, the estimate $|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|\leq\sqrt{|{\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}|}$ also means we lose half of the rate of convergence, a loss that may not actually be observed in practice, noting that the first error bound (23) also holds.
4 Examples
4.1 A parametric diffusion problem
A specific example is the parametric diffusion problem
$\begin{cases}-\nabla\cdot\big{(}a({\boldsymbol{x}},{\boldsymbol{y}})\nabla u({\boldsymbol{x}},{\boldsymbol{y}})\big{)}=f({\boldsymbol{x}}),&{\boldsymbol{x}}\in D,\quad{\boldsymbol{y}}\in Y=[0,1]^{s},\\ u({\boldsymbol{x}},{\boldsymbol{y}})=0,&{\boldsymbol{x}}\in\partial D,\end{cases}$ (61)
where $D$ is a bounded spatial domain in ${\mathbb{R}}^{d}$ with $d\in\{1,2,3\}$, $f\in L^{2}(D)$, and the diffusion coefficient $a({\boldsymbol{x}},{\boldsymbol{y}})$ is a given positive function of ${\boldsymbol{x}}$ and ${\boldsymbol{y}}$. We consider two models:
  • (a)
    Affine (non-periodic) model (see e.g., [7])
    $a({\boldsymbol{x}},{\boldsymbol{y}}):=\psi_{0}({\boldsymbol{x}})+\sum_{j=1}^{s}(y_{j}-\tfrac{1}{2})\,\psi_{j}({\boldsymbol{x}}).$ (62)
  • (b)
    Periodic model (see [23, 24] but without the $1/\sqrt{6}$ scaling)
    $a({\boldsymbol{x}},{\boldsymbol{y}}):=\psi_{0}({\boldsymbol{x}})+\sum_{j=1}^{s}\sin(2\pi y_{j})\,\psi_{j}({\boldsymbol{x}}).$ (63)
In both cases we assume that there exist constants $a_{\min}$ and $a_{\max}$ (different for the affine and periodic models) such that $0<a_{\min}\leq a({\boldsymbol{x}},{\boldsymbol{y}})\leq a_{\max}<\infty$, and that the functions $\psi_{j}\in L^{\infty}(D)$ satisfy additional assumptions. An important assumption when the dimensionality $s$ is high (or infinite) is that there exists a “summability exponent” $p^{*}\in(0,1)$ such that $\sum_{j\geq 1}\|\psi_{j}\|_{\infty}^{p^{*}}<\infty$.
As a concrete example, we consider a simple observable which is a linear functional on the PDE solution:
$G({\boldsymbol{y}}):={\mathcal{G}}(u(\cdot,{\boldsymbol{y}}))=\int_{D^{\prime}}u({\boldsymbol{x}},{\boldsymbol{y}})\,{\mathrm{d}}{\boldsymbol{x}},$ (64)
where $D^{\prime}$ may be one half/quadrant/octant of the domain $D$. Thus $N_{\mathrm{obs}}=1$. More generally, the observables might be the vector of values of the solution $u$ at the points of a finite element mesh, in which case $N_{\mathrm{obs}}$ could be the number of interior finite element mesh points.
The space $V=H^{1}_{0}(D)$ is the typical Sobolev space with first order derivatives with respect to ${\boldsymbol{x}}\in D$, with zero boundary condition and norm $\|v\|_{V}:=\|\nabla v\|_{L^{2}(D)}$. The space $V^{\prime}=H^{-1}(D)$ is its dual. The parametric regularity of the PDE solutions are known in both cases:
  • (a)
    For the affine (non-periodic) model (62) we have (see e.g., [31])
    $\|\partial^{\boldsymbol{\nu}}u(\cdot,{\boldsymbol{y}})\|_{V}\leq\frac{\|f\|_{V^{\prime}}}{a_{\min}}\,|{\boldsymbol{\nu}}|!\,{\boldsymbol{b}}^{\boldsymbol{\nu}},\qquad b_{j}:=\frac{\|\psi_{j}\|_{L_{\infty}(D)}}{a_{\min}}.$ (65)
    The target function (64) satisfies the regularity bound (40) with $b_{j}$ as in (65) and
    $C:=\frac{\|{\mathcal{G}}\|_{V^{\prime}}\,\|f\|_{V^{\prime}}}{a_{\min}}.$ (66)
  • (b)
    For the periodic model (63) we have (see [24])
    $\|\partial^{\boldsymbol{\nu}}u(\cdot,{\boldsymbol{y}})\|_{V}\leq\frac{\|f\|_{V^{\prime}}}{a_{\min}}(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}|{\boldsymbol{m}}|!\,{\boldsymbol{b}}^{\boldsymbol{m}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}),$ (67)
    where $b_{j}$ is defined as in (65), but with a different $a_{\min}$ appropriate for the periodic model. The target function (64) satisfies the regularity bound (45) with $b_{j}$ as in (65) and $C$ as in (66), but again with a different $a_{\min}$. Note that we do not have a $1/\sqrt{6}$ factor in (63) unlike in [24, 23].
4.2 An algebraic equation
We consider an algebraic equation $a({\boldsymbol{y}})\,G({\boldsymbol{y}})=1$, ${\boldsymbol{y}}\in U_{s}$, which mimics the features of the parametric PDE (61) but avoids the spatial variable ${\boldsymbol{x}}$ and the added complexity of a finite element solver. The target function we seek to approximate by a DNN is
$G({\boldsymbol{y}}):=\frac{1}{a({\boldsymbol{y}})},$ (68)
where we consider two models for the denominator (compare to (62) and (63)):
  • (a)
    Affine (non-periodic) algebraic model
    $a({\boldsymbol{y}}):=1+\sum_{j=1}^{s}(y_{j}-\tfrac{1}{2})\,\psi_{j},\qquad\psi_{j}:=\frac{\eta}{j^{q}}\mbox{ for }q>1,\qquad a_{\min}:=1-\frac{\eta\,\zeta(q)}{2},$ (69)
    where $\zeta(\cdot)$ is again the Riemann zeta function. The target function (68) satisfies the regularity bound (40) with $b_{j}=\psi_{j}/a_{\min}$ and $C=1/a_{\min}$.
  • (b)
    Periodic algebraic model
    $a({\boldsymbol{y}}):=1+\sum_{j=1}^{s}\sin(2\pi y_{j})\,\psi_{j},\qquad\psi_{j}:=\frac{\eta}{j^{q}}\mbox{ for }q>1,\qquad a_{\min}:=1-\eta\,\zeta(q).$ (70)
    The target function (68) satisfies the regularity bound (45) with $b_{j}=\psi_{j}/a_{\min}$ and $C=1/a_{\min}$, remembering that $a_{\min}$ is different between the affine and periodic cases.
5 Numerical experiments
In our experiments we will train DNNs to approximate the target function (68) from the periodic algebraic model (70). We aim to demonstrate that restricting network parameters to match the regularity features of the target function as in Theorem 4 can improve the resulting DNN.
5.1 Tailored regularization
We train our DNN to obtain
$\theta^{*}:=\operatorname*{\mathrm{arg\,min}}_{\theta\in\Theta}\bigg{(}\frac{1}{N}\sum_{k=1}^{N}\big{\|}G({\boldsymbol{y}}_{k})-G_{\theta}^{[L]}({\boldsymbol{y}}_{k})\big{\|}_{2}^{2}+\lambda\,\|\theta\|_{2}^{2}+\lambda_{1}\,{\mathcal{R}}_{1}(\theta)\bigg{)}.$
The case $\lambda_{1}=0$ corresponds to the standard $\ell_{2}$ regularization. Otherwise we have our tailored regularization designed based on Theorem 4:
${\mathcal{R}}_{1}(\theta)$ $:=\frac{1}{s}\sum_{j=1}^{s}\frac{1}{d_{1}}\sum_{p=1}^{d_{1}}\Big{(}W_{0,p,j}^{2}\,\frac{L^{2}}{b_{j}^{2}}\Big{)}^{m/2},$ (71)
with a positive even integer $m=6$.
The regularization term ${\mathcal{R}}_{1}(\theta)$ is chosen to encourage the condition (38) from Theorem 4, i.e., $\|W_{0,:,j}\|_{\infty}\leq\beta_{j}\leq b_{j}/S_{L}$ for all $j=1,\ldots,s$, which is equivalent to
$\max_{1\leq j\leq s}\max_{1\leq p\leq d_{1}}|W_{0,p,j}|\,\frac{S_{L}}{b_{j}}\leq 1.$
In an ideal scenario we could have $\xi=\tau=1$ in (15) and $\rho=1$ in (36), leading to $S_{L}\leq L$ in (37). While this is unlikely to be true in general, for simplicity we replace $S_{L}$ by $L$ in designing our regularization term, to at least capture the fact that $S_{L}$ will generally increase with increasing $L$. Ultimately this only affects the scaling of the regularization term.
Recall that $\max_{1\leq i\leq n}|a_{i}|=\lim_{m\to\infty}(\frac{1}{n}\sum_{i=1}^{n}|a_{i}|^{m})^{1/m}$. Since we need ${\mathcal{R}}_{1}(\theta)$ to be smooth so that an automatic differentiation routine can compute $\nabla_{\theta}{\mathcal{R}}_{1}(\theta)$, we replace the two maximums by the average of the terms raised to an even power $m=6$ as follows:
$\bigg{[}\frac{1}{s}\sum_{j=1}^{s}\bigg{(}\bigg{[}\frac{1}{d_{1}}\sum_{p=1}^{d_{1}}\bigg{(}|W_{0,p,j}|\,\frac{L}{b_{j}}\bigg{)}^{m}\bigg{]}^{1/m}\bigg{)}^{m}\bigg{]}^{1/m}\leq 1,$
which simplifies to
$\frac{1}{s}\sum_{j=1}^{s}\frac{1}{d_{1}}\sum_{p=1}^{d_{1}}\bigg{(}W_{0,p,j}^{2}\,\frac{L^{2}}{b_{j}^{2}}\bigg{)}^{m/2}\leq 1^{m}=1.$
The even power allows us to omit the absolute values. We note that the function $x^{6}$ stays mostly constant and below $1$ for $x\leq 1$ and grows very quickly for $x>1$.
The conditions (36) and (39) restrict the magnitude of $R_{1},\ldots,R_{L-1}$ and $R_{L}$. These are already controlled by the standard $\ell_{2}$ regularization term $\lambda\,\|\theta\|_{2}^{2}$. The condition (38) is the most important of the three conditions, since the rate of decay of the sequence $\beta_{j}$ will affect the empirical convergence rate for QMC methods and has a direct impact on whether or not the implied constant grows exponentially with $s$.
After training, from our trained DNN we can compute $\beta_{1},\ldots,\beta_{s}$, and thus obtain a value of $\kappa$ satisfying (60). We can also compute $R_{1},\ldots,R_{L-1}$, thus obtain a value of $\rho$ satisfying (36). Furthermore, we can compute $R_{L}$, $C_{L}$ and $S_{L}$ defined in (18). We see from (37) that $S_{L}$ will generally scale like $\rho^{L}$.
In practice, if $C_{L}\leq C$ is not true as demanded in (39), then we can adjust our norm bounds in Theorem 4 with $C$ replaced by $C_{L}$. If $\kappa\leq 1/S_{L}$ is not true, i.e., $\beta_{j}\leq b_{j}/S_{L}$ for all $j$ is not true as demanded in (38) (and most likely it will not be true), then we can adjust our norm bounds in Theorem 4 with each $b_{j}$ replaced by $\kappa\,S_{L}\,b_{j}$. Consequently, as indicated at the end of Theorem 5, the generalization error bounds still hold, but with enlarged implied constants dependent on the value of $\kappa\,S_{L}$.
5.2 Training with full-batch Adam algorithm
We utilize the Python-based open-source deep learning framework PyTorch [43] to execute a full-batch Adam (adaptive moment estimation) optimization algorithm [29] which is a gradient descent method. The algorithm begins with a specified initial parameter configuration (random “Glorot initialization”) and employs a “learning rate” (step size) of $10^{-4}$. The optimization process halts either when the training error ${\mathcal{E}}_{T}$ (see (21)) reaches the threshold ${\tt tol}=10^{-3}$, or upon completing a maximum number of “epochs” (steps) of $40000$.
We use a randomly shifted version of a lattice point set as our training points $\{{\boldsymbol{y}}_{k}\}_{k\geq 1}$, and progressively train the network using $N=2^{5},\ldots,2^{12}$ points. For simplicity, we take one “off-the-shelf” embedded lattice rule generating vector, constructed according to the method described in [9].
We compare the training error ${\mathcal{E}}_{T}$ to an estimate of the generalization error ${\mathcal{E}}_{G}$ (see (20)) given by
${\mathcal{E}}_{G}\approx\widetilde{{\mathcal{E}}}_{G}:=\bigg{(}\frac{1}{M}\sum_{i=1}^{M}\big{(}G({\boldsymbol{t}}_{i})-G_{\theta}^{[L]}({\boldsymbol{t}}_{i})\big{)}^{2}\bigg{)}^{1/2},$
where we use one different and independent random shift with the same embedded lattice rule with a much higher number of points $M=2^{15}\gg N$. We then compute an estimate of the generalization gap by |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$. Recall that $|{\mathcal{E}}_{G}^{2}-{\mathcal{E}}_{T}^{2}|$ is the quadrature error for the integrand $(G-G_{\theta}^{L})^{2}$, so we expect to observe a bound of the form
${\mathcal{E}}_{G}\leq{\mathcal{E}}_{T}+|{\mathcal{E}}_{G}-{\mathcal{E}}_{T}|\leq{\mathcal{E}}_{T}+\sqrt{e^{\rm wor}_{N}({\boldsymbol{z}})\,\|(G-G_{\theta}^{L})^{2}\|_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}}\leq 10^{-3}+{\mathcal{O}}\big{(}N^{-r/2}\big{)},$
with $r$ depending on the three settings in Theorem 5. Ideally, we hope to observe that the estimated generalization error ~$\widetilde{{\mathcal{E}}}_{G}$ is consistently reduced when tailored regularization is applied. Additionally, we hope to see a higher convergence rate for the generalization gap |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ when tailored regularization is employed and QMC points are used instead of standard Monte Carlo (MC) points.
Figure 1: Values of ${\mathcal{E}}_{T}$ ($\circ$ green circles), ~$\widetilde{{\mathcal{E}}}_{G}$ ($\bullet$ red dots), and |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ ( $\blacksquare$ blue squares) as $N$ increases, with $L=3$, $N_{\rm obs}=1$, $d_{\ell}=32$, $s=50$.
Figure 2: Values of ${\mathcal{E}}_{T}$ ($\circ$ green circles), ~$\widetilde{{\mathcal{E}}}_{G}$ ($\bullet$ red dots), and |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ ( $\blacksquare$ blue squares) as $N$ increases, with $L=12$, $N_{\rm obs}=1$, $d_{\ell}=30$, $s=50$.
For the numerical experiments, we will investigate the periodic algebraic model (68) and (70) with
$\eta=0.5,\quad q=2.5,$
and we will train a periodic DNN (4) with the sigmoid activation function using two sets of hyperparameters:
  1. 1.
    $L=3$,   $N_{\rm obs}=1=d_{L+1}=d_{4}$,    $d_{3}=d_{2}=d_{1}=32$, $d_{0}=s=50$  ($3777$ parameters);
  2. 2.
    $L=12$, $N_{\rm obs}=1=d_{L+1}=d_{13}$, $d_{12}=\cdots=d_{1}=30$, $d_{0}=s=50$  ($11791$ parameters);
that is, one relatively small network and one larger network. While some may argue that these networks are small, we point out that “tiny” neural networks have recently been shown to be very successful in computer graphics, see [39, 40, 51].
In Figure 1, for the first set of hyperparameters ($L=3$), we plot the values of the training error ${\mathcal{E}}_{T}$ ($\circ$ green circles), the estimated generalization error ~$\widetilde{{\mathcal{E}}}_{G}$ ($\bullet$ red dots), and the generalization gap |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ ( $\blacksquare$ blue squares) for $N=2^{5},\ldots,2^{12}$, where each of the three quantities is averaged over $20$ different random “Glorot” initializations of the network and random shifts of the lattice rule. The top two graphs (labelled “standard regularization”) were obtained with regularization parameter $\lambda=10^{-8}$ but without our tailored regularization (so $\lambda_{1}=0$), while the bottom two graphs (labelled “tailored regularization”) were obtained with $\lambda=10^{-8}$ and with our tailored regularization $\lambda_{1}=10^{-8}$. The two graphs on the left (labelled “QMC”) were obtained by training with randomly shifted embedded lattice points as mentioned above, while the two graphs on the right (labelled “MC”) were obtained with uniformly distributed random points for comparison. In Figure 2 we give the same plots for the second set of hyperparameters ($L=12$).
We observe in all four cases in Figure 1 ($L=3$) that the training error ${\mathcal{E}}_{T}$ ($\circ$) is able to reach the threshold ${\tt tol}=10^{-3}$ for most values of $N$. With just the standard regularization (i)–(ii), both ~$\widetilde{{\mathcal{E}}}_{G}$ ($\bullet$) and |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ ( $\blacksquare$ ) are large and plateau initially before diving down from $N=2^{11}$, but even at $N=2^{12}$ they are still some distance away from ${\tt tol}=10^{-3}$. On the other hand, with tailored regularization (iii)–(iv), |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ ( $\blacksquare$ ) decays at a rate between ${\mathcal{O}}(N^{-1})$ and ${\mathcal{O}}(N^{-2})$, and quickly crosses below ${\tt tol}=10^{-3}$ by $N=2^{9}$, with ~$\widetilde{{\mathcal{E}}}_{G}$ ($\bullet$) fairly close to ${\tt tol}=10^{-3}$ from $N=2^{10}$.
In Figure 2 ($L=12$) we observe that, with just the standard regularization (i)–(ii), the training error ${\mathcal{E}}_{T}$ ($\circ$) cannot reach the threshold ${\tt tol}=10^{-3}$ until $N=2^{7}$, and both ~$\widetilde{{\mathcal{E}}}_{G}$ ($\bullet$) and |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ ( $\blacksquare$ ) stay some distance away from ${\tt tol}=10^{-3}$ even at $N=2^{12}$. On the other hand, with tailored regularization (iii)–(iv), the training error ${\mathcal{E}}_{T}$ ($\circ$) cannot reach the threshold ${\tt tol}=10^{-3}$ for all values of $N$. While this may be disappointing, we observe that ~$\widetilde{{\mathcal{E}}}_{G}$ ($\bullet$) follows ${\mathcal{E}}_{T}$ ($\circ$) closely, and the gap |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ ( $\blacksquare$ ) crosses below ${\mathcal{E}}_{T}$ ($\circ$) from $N=2^{7}$ and decays at a rate between ${\mathcal{O}}(N^{-1})$ and ${\mathcal{O}}(N^{-2})$ as before. Note that the learning rate is constant. We expect the results to improve once we introduce a strategy to reduce the learning rate during training.
In summary, tailored regularization provides consistent improvement in all our experiments, and it benefits QMC as well as MC. A surprise is that MC performs just as well as QMC, with QMC perhaps having a slightly better empirical convergence rate for the gap |~$|\widetilde{{\mathcal{E}}}_{G}-{\mathcal{E}}_{T}|$ ( $\blacksquare$ ). The theory in our paper does not directly apply to MC, although loosely speaking we may replace the worst case error bound $e^{\rm wor}_{N}({\boldsymbol{z}})\,\|(G-G_{\theta}^{L})^{2}\|_{{\mathcal{W}}_{\alpha,{\boldsymbol{\gamma}}}}$ by the variance of $(G-G_{\theta}^{L})^{2}$ divided by $\sqrt{N}$. A possible explanation for the success of MC with tailored regularization may be that tailored regularization reduces the variance of the squared difference.
In Figure 3, for one random initialization and one random shift with the first set of hyperparameters ($L=3$) and $N=2^{5}$, we plot the values of $\log(\beta_{j})$ as blue dots and $\log(b_{j}/L)$ as a black line for $j=1,\ldots,s$. We see that the sequence $(\beta_{j})$ with tailored regularization (iii)–(iv) decays closely following the sequence $(b_{j}/L)$ but there is a scaling constant. In fact, for this instance of case (iii) we achieve $\beta_{j}\leq\kappa\,b_{j}$ for all $j=1,\ldots,s$ with $\kappa=1.4851$ and for the matrices $W_{1},\ldots,W_{L-1}$ we achieve $R_{\ell}\leq\rho$ for all $\ell=1,\ldots,L-1$ with $\rho=18.371$. Results for the second set of hyperparameters ($L=12$) are shown in Figure 4, exhibiting a similar trend, where for this instance of case (iii) we achieve $\kappa=1.2951$ and $\rho=10.997$.
Figure 3: Values of $\log(\beta_{j})$ ($\bullet$ blue dots) and $\log(b_{j}/L)$ (black line) for $j=1,\ldots,s$ for one random initialization and one random shift, with $L=3$, $N_{\rm obs}=1$, $d_{\ell}=32$, $s=50$, $N=2^{5}$.
Figure 4: Values of $\log(\beta_{j})$ ($\bullet$ blue dots) and $\log(b_{j}/L)$ (black line) for $j=1,\ldots,s$, for one random initialization and one random shift, with $L=12$, $N_{\rm obs}\!=\!1$, $d_{\ell}\!=\!30$, $s=50$, $N=2^{5}$.
6 Technical proofs
6.1 Proof of regularity bound for non-periodic DNN
Recall the notation from Section 2, including ${\mathbb{N}}_{0}:=\{0,1,2,\ldots\}$ and :=$\mathcal{I}:=\{{\boldsymbol{\nu}}\in{\mathbb{N}}_{0}^{\infty}:|{\boldsymbol{\nu}}|<\infty\}$. In the following we will use the Kronecker delta function: $\delta_{{\boldsymbol{\nu}},{\boldsymbol{0}}}$ is $1$ if ${\boldsymbol{\nu}}={\boldsymbol{0}}$ and is 00 otherwise. Let ${\boldsymbol{e}}_{j}$ denote the multiindex whose $j$th component is $1$ and all other components are 00.
First we prove a generic technical result involving a sharp recursion. The key point is that this result separates the “product” (see $\beta_{j}$ below) contributions from other “order dependent” (see $\Gamma_{n}$ below) contributions of the sequence.
Lemma 6
Let the nonnegative sequence $({\mathbb{A}}_{{\boldsymbol{\nu}},\lambda})_{{\boldsymbol{\nu}}\in\mathcal{I},\,\lambda\in{\mathbb{N}}_{0}}$ satisfy ${\mathbb{A}}_{{\boldsymbol{\nu}},0}=\delta_{{\boldsymbol{\nu}},{\boldsymbol{0}}}$, ${\mathbb{A}}_{{\boldsymbol{\nu}},\lambda}=0$ for $\lambda>|{\boldsymbol{\nu}}|$, and otherwise satisfy the recursion
[…]
for some positive $R$, $(\beta_{j})_{j\geq 1}$ and $(\Gamma_{n})_{n\geq 0}$. Then for all ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ and $1\leq\lambda\leq|{\boldsymbol{\nu}}|$ we have
${\mathbb{A}}_{{\boldsymbol{\nu}},\lambda}=R^{\lambda}\,{\boldsymbol{\beta}}^{\boldsymbol{\nu}}\,{\mathbb{B}}_{|{\boldsymbol{\nu}}|,\lambda}\,,$ (73)
with
${\mathbb{B}}_{n,1}:=\Gamma_{n}\quad\mbox{for $n\geq 1$},\qquad{\mathbb{B}}_{n,\lambda}:=\sum_{i=\lambda-1}^{n-1}\binom{n-1}{i}\Gamma_{n-i}\,{\mathbb{B}}_{i,\lambda-1}\quad\mbox{for $n\geq\lambda\geq 2$}.$ (74)
If (72) is replaced by an inequality $\leq$ then so is (73).
Proof.
We prove (73) by induction on $\lambda$. Consider first the case $\lambda=1$. Then for all ${\boldsymbol{\nu}}\in\mathcal{I}$ and all $j\geq 1$ we have from (72) (since only the ${\boldsymbol{m}}={\boldsymbol{0}}$ term remains)
${\mathbb{A}}_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},1}=R\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}\,\Gamma_{|{\boldsymbol{\nu}}|+1}=R\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}\,{\mathbb{B}}_{|{\boldsymbol{\nu}}|+1,1},$
where we used $\Gamma_{|{\boldsymbol{\nu}}|+1}={\mathbb{B}}_{|{\boldsymbol{\nu}}|+1,1}$ from (74). This proves that (73) holds for $\lambda=1$.
For $\lambda\geq 2$, we assume as induction hypothesis that (c.f. (73))
${\mathbb{A}}_{{\boldsymbol{m}},\lambda-1}=R^{\lambda-1}\,{\boldsymbol{\beta}}^{\boldsymbol{m}}\,{\mathbb{B}}_{|{\boldsymbol{m}}|,\lambda-1}$
for all ${\boldsymbol{m}}\neq{\boldsymbol{0}}$ with $|{\boldsymbol{m}}|\geq\lambda-1$. Recall that ${\mathbb{A}}_{{\boldsymbol{m}},\lambda-1}=0$ for $|{\boldsymbol{m}}|<\lambda-1$. Then for all ${\boldsymbol{\nu}}\in\mathcal{I}$ and all $j\geq 1$ we have from (72) that
${\mathbb{A}}_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}$ $=R\sum_{\begin{subarray}{c}{\boldsymbol{m}}\leq{\boldsymbol{\nu}}\\ |{\boldsymbol{m}}|\geq\lambda-1\end{subarray}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}\,\Gamma_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|+1}\,R^{\lambda-1}\,{\boldsymbol{\beta}}^{\boldsymbol{m}}\,{\mathbb{B}}_{|{\boldsymbol{m}}|,\lambda-1}$
$=R^{\lambda}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}\sum_{\begin{subarray}{c}{\boldsymbol{m}}\leq{\boldsymbol{\nu}}\\ |{\boldsymbol{m}}|\geq\lambda-1\end{subarray}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\Gamma_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|+1}\,{\mathbb{B}}_{|{\boldsymbol{m}}|,\lambda-1}\,$
$=R^{\lambda}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}\sum_{i=\lambda-1}^{|{\boldsymbol{\nu}}|}\sum_{\begin{subarray}{c}{\boldsymbol{m}}\leq{\boldsymbol{\nu}}\\ |{\boldsymbol{m}}|=i\end{subarray}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\Gamma_{|{\boldsymbol{\nu}}|-i+1}\,{\mathbb{B}}_{i,\lambda-1}=R^{\lambda}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}\underbrace{\sum_{i=\lambda-1}^{|{\boldsymbol{\nu}}|}\binom{|{\boldsymbol{\nu}}|}{i}\Gamma_{|{\boldsymbol{\nu}}|-i+1}\,{\mathbb{B}}_{i,\lambda-1}}_{=\,{\mathbb{B}}_{|{\boldsymbol{\nu}}|+1,\lambda}},$
where we used the combinatorial identity (e.g., consider the number of ways to select $i$ distinct balls from some baskets containing a total number of $|{\boldsymbol{\nu}}|$ distinct balls)
$\sum_{\begin{subarray}{c}{\boldsymbol{m}}\leq{\boldsymbol{\nu}}\\ |{\boldsymbol{m}}|=i\end{subarray}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}=\binom{|{\boldsymbol{\nu}}|}{i},$ (75)
together with the recursive definition (74). This completes the induction proof.
If (72) is replaced by an inequality $\leq$ , then it is easy to see from the proof that we can replace the induction hypothesis also by an inequality $\leq$ and this will lead to an inequality $\leq$ in the final result (73). ∎
  • Proof of Theorem 1(a).
    Consider first a shallow network with one hidden layer
    $G_{\theta}^{[1]}({\boldsymbol{y}})=W_{1}\,\sigma(W_{0}\,{\boldsymbol{y}}+{\boldsymbol{v}}_{0})+{\boldsymbol{v}}_{1}=\bigg{[}\sum_{q=1}^{d_{1}}W_{1,p,q}\,\sigma\bigg{(}\sum_{r=1}^{d_{0}}W_{0,q,r}\,y_{r}+v_{0,q}\bigg{)}+v_{1,p}\bigg{]}_{p=1}^{d_{2}},$
    where $d_{0}=s$. For any ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ and any index $p$, it is easy to see that by the chain rule every derivative with respect to a variable $y_{j}$ will yield an extra constant factor $W_{0,q,j}$, leading to
    $\partial^{\boldsymbol{\nu}}G_{\theta}^{[1]}({\boldsymbol{y}})_{p}=\sum_{q=1}^{d_{1}}W_{1,p,q}\,\sigma^{(|{\boldsymbol{\nu}}|)}\bigg{(}\sum_{r=1}^{d_{0}}W_{0,q,r}\,y_{r}+v_{0,q}\bigg{)}\,\prod_{j=1}^{d_{0}}W_{0,q,j}^{\nu_{j}},$
    and thus
    $|\partial^{\boldsymbol{\nu}}G_{\theta}^{[1]}({\boldsymbol{y}})_{p}|\leq\sum_{q=1}^{d_{1}}|W_{1,p,q}|\,A_{|{\boldsymbol{\nu}}|}\prod_{j=1}^{d_{0}}|W_{0,q,j}|^{\nu_{j}}\leq R_{1}\,A_{|{\boldsymbol{\nu}}|}\prod_{j=1}^{s}\beta_{j}^{\nu_{j}}=R_{1}\,{\boldsymbol{\beta}}^{\boldsymbol{\nu}}\,A_{|{\boldsymbol{\nu}}|}=R_{1}\,{\boldsymbol{\beta}}^{\boldsymbol{\nu}}\,\Gamma_{|{\boldsymbol{\nu}}|}^{[1]}\,,$
    where we used the assumptions (7), (8), (9), together with $A_{|{\boldsymbol{\nu}}|}=\Gamma_{|{\boldsymbol{\nu}}|}^{[1]}$ from (12). Hence (10) holds for $\ell=1$.
    Now consider $\ell\geq 2$ and suppose the result (10) holds for depth $\ell-1$. We can write
    $G_{\theta}^{[\ell]}({\boldsymbol{y}})=W_{\ell}\,\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}}))+{\boldsymbol{v}}_{\ell}=\bigg{[}\sum_{q=1}^{d_{\ell}}W_{\ell,p,q}\,\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q})+v_{\ell,p}\bigg{]}_{p=1}^{d_{\ell+1}}.$
    For any ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ and any index $p$, we have
    $\partial^{\boldsymbol{\nu}}G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}$ $=\sum_{q=1}^{d_{\ell}}W_{\ell,p,q}\,\partial^{\boldsymbol{\nu}}\big{(}\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q})\big{)}.$ (76)
    For each index $q$, the recursive form of the multivariate Faà di Bruno formula gives (see [45, formula (3.1) and (3.5)] with the dimensionality $m$ there set to $1$)
    $\partial^{\boldsymbol{\nu}}\big{(}\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q})\big{)}$ $=\sum_{\lambda=1}^{|{\boldsymbol{\nu}}|}\sigma^{(\lambda)}(G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q})\,\alpha_{{\boldsymbol{\nu}},\lambda}^{[\ell-1,q]}({\boldsymbol{y}}),$ (77)
    where the auxiliary functions $\alpha_{{\boldsymbol{\nu}},\lambda}^{[\ell-1,q]}({\boldsymbol{y}})$ are defined recursively by $\alpha_{{\boldsymbol{\nu}},0}^{[\ell-1,q]}({\boldsymbol{y}}):=\delta_{{\boldsymbol{\nu}},{\boldsymbol{0}}}$, $\alpha_{{\boldsymbol{\nu}},\lambda}^{[\ell-1,q]}({\boldsymbol{y}}):=0$ for $\lambda>|{\boldsymbol{\nu}}|$, and otherwise
    $\alpha_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}^{[\ell-1,q]}({\boldsymbol{y}}):=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\partial^{{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q}\,\alpha_{{\boldsymbol{m}},\lambda-1}^{[\ell-1,q]}({\boldsymbol{y}}).$
    Using the induction hypothesis (10) for depth $\ell-1$, we obtain
    $|\alpha_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}^{[\ell-1,q]}({\boldsymbol{y}})|$ $\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}|\partial^{{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q}|\,|\alpha_{{\boldsymbol{m}},\lambda-1}^{[\ell-1,q]}({\boldsymbol{y}})|$ (78)
    $\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}R_{\ell-1}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}\,\Gamma_{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|+1}^{[\ell-1]}\,|\alpha_{{\boldsymbol{m}},\lambda-1}^{[\ell-1,q]}({\boldsymbol{y}})|.$
    Lemma 6 (with inequalities $\leq$) then yields for all ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$,
    $|\alpha_{{\boldsymbol{\nu}},\lambda}^{[\ell-1,q]}({\boldsymbol{y}})|\leq R_{\ell-1}^{\lambda}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\,{\mathbb{B}}_{|{\boldsymbol{\nu}}|,\lambda}=R_{\ell-1}^{\lambda}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\,{\mathbb{B}}_{|{\boldsymbol{\nu}}|,\lambda}^{[\ell-1]},$ (79)
    where the sequence ${\mathbb{B}}_{n,\lambda}$ is defined by (74) with $\Gamma_{n}=\Gamma_{n}^{[\ell-1]}$, and this sequence ${\mathbb{B}}_{n,\lambda}$ is precisely ${\mathbb{B}}_{n,\lambda}^{[\ell-1]}$ defined in (13). Since the same upper bound (79) holds for all indices $q$, we conclude from (76) and (77) that
    $|\partial^{\boldsymbol{\nu}}G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}|$ $\leq\sum_{q=1}^{d_{\ell}}|W_{\ell,p,q}|\sum_{\lambda=1}^{|{\boldsymbol{\nu}}|}A_{\lambda}\,R_{\ell-1}^{\lambda}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\,{\mathbb{B}}_{|{\boldsymbol{\nu}}|,\lambda}^{[\ell-1]}\leq R_{\ell}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\sum_{\lambda=1}^{|{\boldsymbol{\nu}}|}A_{\lambda}\,R_{\ell-1}^{\lambda}\,{\mathbb{B}}_{|{\boldsymbol{\nu}}|,\lambda}^{[\ell-1]}=R_{\ell}\,{\boldsymbol{\beta}}^{{\boldsymbol{\nu}}}\,\Gamma_{|{\boldsymbol{\nu}}|}^{[\ell]},$
[…]
The bound in Theorem 1(a) is reasonably tight, with inequalities arising only from bounding the matrix elements (7) and (8), the derivatives of the activation function (9), and from taking absolute values in (78). Otherwise the induction argument from Lemma 6 is sharp.
6.2 Proof of regularity bound for periodic DNN
As in the non-periodic case, we prove here a generic technical result involving a sharp recursion for the periodic case. The key point is that this result separates the “product” contributions from other “order dependent” contributions of the sequence, as well as the dependence on Stirling numbers of the second kind.
Lemma 7
Let the nonnegative sequence $({\mathbb{A}}_{{\boldsymbol{\nu}},\lambda})_{{\boldsymbol{\nu}}\in\mathcal{I},\,\lambda\in{\mathbb{N}}_{0}}$ satisfy ${\mathbb{A}}_{{\boldsymbol{\nu}},0}=\delta_{{\boldsymbol{\nu}},{\boldsymbol{0}}}$, ${\mathbb{A}}_{{\boldsymbol{\nu}},\lambda}=0$ for $\lambda>|{\boldsymbol{\nu}}|$, and otherwise satisfy the recursion
${\mathbb{A}}_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}=R\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\bigg{(}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}{\boldsymbol{\beta}}^{{\boldsymbol{w}}}\,\Gamma_{|{\boldsymbol{w}}|}\,$ ${\mathcal{S}}({\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}})\bigg{)}{\mathbb{A}}_{{\boldsymbol{m}},\lambda-1},$
[…]
for some positive $R$, $(\beta_{j})_{j\geq 1}$ and $(\Gamma_{n})_{n\geq 0}$. Then for all ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ and $1\leq\lambda\leq|{\boldsymbol{\nu}}|$ we have
${\mathbb{A}}_{{\boldsymbol{\nu}},\lambda}=R^{\lambda}\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{u}}\leq{\boldsymbol{\nu}}}}}{{\scriptstyle{|{\boldsymbol{u}}|\geq\lambda}}}}{\boldsymbol{\beta}}^{\boldsymbol{u}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{u}}),$ (81)
with ${\mathbb{B}}_{n,\lambda}$ defined as in (74) but with $\Gamma_{n}$ as used in (7) rather than in (72). If (7) is replaced by an inequality $\leq$ then so is (81).
Proof.
We prove (81) by induction on $\lambda$. Consider first the case $\lambda=1$. Then for all ${\boldsymbol{\nu}}\in\mathcal{I}$ and all $j\geq 1$ we have from (7) (since only the ${\boldsymbol{m}}={\boldsymbol{0}}$ term remains)
${\mathbb{A}}_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},1}=R\sum_{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}{\boldsymbol{\beta}}^{\boldsymbol{w}}\,\Gamma_{|{\boldsymbol{w}}|}\,{\mathcal{S}}({\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}})=R\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}}}{{\scriptstyle{|{\boldsymbol{w}}|\geq 1}}}}{\boldsymbol{\beta}}^{\boldsymbol{w}}\,{\mathbb{B}}_{|{\boldsymbol{w}}|,1}\,{\mathcal{S}}({\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}}),$
which agrees with (81) since $\Gamma_{|{\boldsymbol{w}}|}={\mathbb{B}}_{|{\boldsymbol{w}}|,1}$ as defined in (74) and ${\mathcal{S}}({\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},{\boldsymbol{0}})=0$. This proves that (81) holds for $\lambda=1$.
For $\lambda\geq 2$, we assume as induction hypothesis that (c.f. (81))
${\mathbb{A}}_{{\boldsymbol{m}},\lambda-1}=R^{\lambda-1}\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{u}}\leq{\boldsymbol{m}}}}}{{\scriptstyle{|{\boldsymbol{u}}|\geq\lambda-1}}}}{\boldsymbol{\beta}}^{\boldsymbol{u}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda-1}\,{\mathcal{S}}({\boldsymbol{m}},{\boldsymbol{u}})$
for all ${\boldsymbol{m}}\neq{\boldsymbol{0}}$ with $|{\boldsymbol{m}}|\geq\lambda-1$. Recall that ${\mathbb{A}}_{{\boldsymbol{m}},\lambda-1}=0$ for $|{\boldsymbol{m}}|<\lambda-1$. Then for all ${\boldsymbol{\nu}}\in\mathcal{I}$ and all $j\geq 1$ we have from (7) that
${\mathbb{A}}_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}$
$=R\sum_{\begin{subarray}{c}{\boldsymbol{m}}\leq{\boldsymbol{\nu}}\\ |{\boldsymbol{m}}|\geq\lambda-1\end{subarray}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\bigg{(}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}{\boldsymbol{\beta}}^{{\boldsymbol{w}}}\,\Gamma_{|{\boldsymbol{w}}|}\,{\mathcal{S}}({\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}})\bigg{)}\,R^{\lambda-1}\sum_{\begin{subarray}{c}{\boldsymbol{u}}\leq{\boldsymbol{m}}\\ |{\boldsymbol{u}}|\geq\lambda-1\end{subarray}}{\boldsymbol{\beta}}^{\boldsymbol{u}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda-1}\,{\mathcal{S}}({\boldsymbol{m}},{\boldsymbol{u}}).$
Next we swap the order of summation twice
${\mathbb{A}}_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}$
$=R^{\lambda}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}{\boldsymbol{\beta}}^{\boldsymbol{w}}\,\Gamma_{|{\boldsymbol{w}}|}\sum_{\begin{subarray}{c}{\boldsymbol{m}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}-{\boldsymbol{w}}\\ |{\boldsymbol{m}}|\geq\lambda-1\end{subarray}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}{\mathcal{S}}({\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}})\,\sum_{\begin{subarray}{c}{\boldsymbol{u}}\leq{\boldsymbol{m}}\\ |{\boldsymbol{u}}|\geq\lambda-1\end{subarray}}{\boldsymbol{\beta}}^{\boldsymbol{u}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda-1}\,{\mathcal{S}}({\boldsymbol{m}},{\boldsymbol{u}})$
$=R^{\lambda}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}{\boldsymbol{\beta}}^{\boldsymbol{w}}\,\Gamma_{|{\boldsymbol{w}}|}\sum_{\begin{subarray}{c}{\boldsymbol{u}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}-{\boldsymbol{w}}\\ |{\boldsymbol{u}}|\geq\lambda-1\end{subarray}}{\boldsymbol{\beta}}^{\boldsymbol{u}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda-1}\sum_{\begin{subarray}{c}{\boldsymbol{u}}\leq{\boldsymbol{m}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}-{\boldsymbol{w}}\\ |{\boldsymbol{m}}|\geq\lambda-1\end{subarray}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}{\mathcal{S}}({\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}})\,{\mathcal{S}}({\boldsymbol{m}},{\boldsymbol{u}}),$ (82)
where we may drop the condition $|{\boldsymbol{m}}|\geq\lambda-1$ because it is already taken care of by the earlier condition $|{\boldsymbol{u}}|\geq\lambda-1$.
The sum over ${\boldsymbol{m}}$ in (6.2) is
$\sum_{{\boldsymbol{u}}\leq{\boldsymbol{m}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}-{\boldsymbol{w}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}{\mathcal{S}}({\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}})\,{\mathcal{S}}({\boldsymbol{m}},{\boldsymbol{u}})=\binom{{\boldsymbol{w}}+{\boldsymbol{u}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{w}}-{\boldsymbol{e}}_{j}}\,{\mathcal{S}}({\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}}+{\boldsymbol{u}}),$ (83)
since (see e.g., [41, formula 26.8.23])
$\sum_{m=u}^{\nu-w}\binom{\nu}{m}{\mathcal{S}}(\nu-m,w)\,{\mathcal{S}}(m,u)=\binom{w+u}{w}{\mathcal{S}}(\nu,w+u),$
while
$\sum_{m=u}^{\nu+1-w}\binom{\nu}{m}{\mathcal{S}}(\nu-m+1,w)\,{\mathcal{S}}(m,u)$
$=\sum_{m=u}^{\nu+1-w}\binom{\nu}{m}\Big{(}w\,{\mathcal{S}}(\nu-m,w)+{\mathcal{S}}(\nu-m,w-1)\Big{)}\,{\mathcal{S}}(m,u)$
$=w\sum_{m=u}^{\nu-w}\binom{\nu}{m}{\mathcal{S}}(\nu-m,w)\,{\mathcal{S}}(m,u)+\sum_{m=u}^{\nu-(w-1)}\binom{\nu}{m}{\mathcal{S}}(\nu-m,w-1)\,{\mathcal{S}}(m,u)$
$=w\binom{w+u}{w}{\mathcal{S}}(\nu,w+u)+\binom{(w-1)+u}{w-1}{\mathcal{S}}(\nu,w-1+u)$
$=\binom{w+u-1}{w-1}\bigg{(}(w+u)\,{\mathcal{S}}(\nu,w+u)+{\mathcal{S}}(\nu,w+u-1)\bigg{)}=\binom{w+u-1}{w-1}{\mathcal{S}}(\nu+1,w+u),$
where we used the recurrence ${\mathcal{S}}(n+1,k)=k\,{\mathcal{S}}(n,k)+{\mathcal{S}}(n,k-1)$ for $0<k<n$ in the first and last equalities. Substituting (83) into (6.2), we arrive at
${\mathbb{A}}_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}$ $=R^{\lambda}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}{\boldsymbol{\beta}}^{\boldsymbol{w}}\,\Gamma_{|{\boldsymbol{w}}|}\sum_{\begin{subarray}{c}{\boldsymbol{u}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}-{\boldsymbol{w}}\\ |{\boldsymbol{u}}|\geq\lambda-1\end{subarray}}{\boldsymbol{\beta}}^{\boldsymbol{u}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda-1}\binom{{\boldsymbol{w}}+{\boldsymbol{u}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{w}}-{\boldsymbol{e}}_{j}}{\mathcal{S}}({\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}}+{\boldsymbol{u}})$
$=R^{\lambda}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}{\boldsymbol{\beta}}^{\boldsymbol{w}}\,\Gamma_{|{\boldsymbol{w}}|}\sum_{\begin{subarray}{c}{\boldsymbol{w}}\leq{\boldsymbol{u}}^{\prime}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}\\ |{\boldsymbol{u}}^{\prime}|\geq|{\boldsymbol{w}}|+\lambda-1\end{subarray}}\beta^{{\boldsymbol{u}}^{\prime}-{\boldsymbol{w}}}\,{\mathbb{B}}_{|{\boldsymbol{u}}^{\prime}-{\boldsymbol{w}}|,\lambda-1}\,\binom{{\boldsymbol{u}}^{\prime}-{\boldsymbol{e}}_{j}}{{\boldsymbol{w}}-{\boldsymbol{e}}_{j}}{\mathcal{S}}({\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},{\boldsymbol{u}}^{\prime})$
$=R^{\lambda}\,\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{u}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}}}{{\scriptstyle{|{\boldsymbol{u}}|\geq\lambda}}}}{\boldsymbol{\beta}}^{\boldsymbol{u}}\bigg{(}\sum_{\begin{subarray}{c}{\boldsymbol{w}}\leq{\boldsymbol{u}}\\ |{\boldsymbol{w}}|\leq|{\boldsymbol{u}}|-(\lambda-1)\end{subarray}}\binom{{\boldsymbol{u}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{w}}-{\boldsymbol{e}}_{j}}\,\Gamma_{|{\boldsymbol{w}}|}\,{\mathbb{B}}_{|{\boldsymbol{u}}-{\boldsymbol{w}}|,\lambda-1}\bigg{)}{\mathcal{S}}({\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},{\boldsymbol{u}}),$
where we substituted ${\boldsymbol{u}}^{\prime}={\boldsymbol{w}}+{\boldsymbol{u}}$ and then swapped the order of summation (and relabeled ${\boldsymbol{u}}^{\prime}$ back to ${\boldsymbol{u}}$ at the same time). We see that implicitly ${\boldsymbol{w}}\geq{\boldsymbol{e}}_{j}$ (otherwise the binomial coefficient will be 00), which leads to the condition $|{\boldsymbol{u}}|\geq\lambda$. We can rewrite the inner sum over ${\boldsymbol{w}}$ as
$\sum_{\begin{subarray}{c}{\boldsymbol{w}}\leq{\boldsymbol{u}}\\ |{\boldsymbol{w}}|\leq|{\boldsymbol{u}}|-(\lambda-1)\end{subarray}}\binom{{\boldsymbol{u}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{w}}-{\boldsymbol{e}}_{j}}\,\Gamma_{|{\boldsymbol{w}}|}\,{\mathbb{B}}_{|{\boldsymbol{u}}-{\boldsymbol{w}}|,\lambda-1}$ $=\sum_{i=1}^{|{\boldsymbol{u}}|-(\lambda-1)}\sum_{\begin{subarray}{c}{\boldsymbol{e}}_{j}\leq{\boldsymbol{w}}\leq{\boldsymbol{u}}\\ |{\boldsymbol{w}}|=i\end{subarray}}\binom{{\boldsymbol{u}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{w}}-{\boldsymbol{e}}_{j}}\,\Gamma_{i}\,{\mathbb{B}}_{|{\boldsymbol{u}}|-i,\lambda-1}$
$=\sum_{i=1}^{|{\boldsymbol{u}}|-(\lambda-1)}\binom{|{\boldsymbol{u}}|-1}{i-1}\,\Gamma_{i}\,{\mathbb{B}}_{|{\boldsymbol{u}}|-i,\lambda-1}$
$=\sum_{i^{\prime}=\lambda-1}^{|{\boldsymbol{u}}|-1}\binom{|{\boldsymbol{u}}|-1}{i^{\prime}}\,\Gamma_{|{\boldsymbol{u}}|-i^{\prime}}\,{\mathbb{B}}_{i^{\prime},\lambda-1}={\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda},$
where we substituted $i^{\prime}=|{\boldsymbol{u}}|-i$ and used the symmetry of the binomial coefficient and the recursive definition of ${\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda}$ in (74). In the second equality above we used (75) to obtain
$\sum_{\begin{subarray}{c}{\boldsymbol{e}}_{j}\leq{\boldsymbol{w}}\leq{\boldsymbol{u}}\\ |{\boldsymbol{w}}|=i\end{subarray}}\binom{{\boldsymbol{u}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{w}}-{\boldsymbol{e}}_{j}}=\sum_{\begin{subarray}{c}{\boldsymbol{w}}^{\prime}\leq{\boldsymbol{u}}-{\boldsymbol{e}}_{j}\\ |{\boldsymbol{w}}^{\prime}|=i-1\end{subarray}}\binom{{\boldsymbol{u}}-{\boldsymbol{e}}_{j}}{{\boldsymbol{w}}^{\prime}}=\binom{|{\boldsymbol{u}}|-1}{i-1}.$
Hence we conclude that
${\mathbb{A}}_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}$ $=R^{\lambda}\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{u}}\leq{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j}}}}{{\scriptstyle{|{\boldsymbol{u}}|\geq\lambda}}}}{\boldsymbol{\beta}}^{\boldsymbol{u}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda}\,{\mathcal{S}}({\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},{\boldsymbol{u}}),$
which proves the induction step for (81). This completes the induction proof.
If (7) is replaced by an inequality $\leq$ , then it is easy to see from the proof that we can replace the induction hypothesis also by an inequality $\leq$ and this will lead to an inequality $\leq$ in the final result (81). ∎
  • Proof of Theorem 1(b).
    The proof for the periodic case is more complicated, hence it is helpful as a preparation step to first consider the case with no hidden layer
    $G_{\theta}^{[0]}({\boldsymbol{y}})=W_{0}\,\sin(2\pi{\boldsymbol{y}})+{\boldsymbol{v}}_{0}=\bigg{[}\sum_{q=1}^{d_{0}}W_{0,p,q}\,\sin(2\pi y_{q})+v_{0,p}\bigg{]}_{p=1}^{d_{1}},$
    where $d_{0}=s$. For ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ and any index $p$, we may differentiate with respect to a single variable $y_{j}$ multiple times but all cross derivatives vanish:
    $\partial^{\boldsymbol{\nu}}G_{\theta}^{[0]}({\boldsymbol{y}})_{p}$ =$=\begin{cases}W_{0,p,j}\,(2\pi)^{k}\sin(2\pi y_{j}+\tfrac{k\pi}{2})&\mbox{if }{\boldsymbol{\nu}}=k{\boldsymbol{e}}_{j},\,k\geq 1,\\ 0&\mbox{otherwise},\end{cases}$ (84)
    from which we conclude that
    $|\partial^{{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}G_{\theta}^{[0]}({\boldsymbol{y}})_{p}|$$\leq\begin{cases}\beta_{j}\,(2\pi)^{k+1}&\mbox{if }{\boldsymbol{m}}=k{\boldsymbol{e}}_{j},\,k\geq 0,\\ 0&\mbox{otherwise},\end{cases}$ (85)
    where we used the assumption (7).
    We are ready to consider a shallow network with one hidden layer
    $G_{\theta}^{[1]}({\boldsymbol{y}})=W_{1}\,\sigma(G_{\theta}^{[0]}({\boldsymbol{y}}))+{\boldsymbol{v}}_{1}=\bigg{[}\sum_{q=1}^{d_{1}}W_{1,p,q}\,\sigma(G_{\theta}^{[0]}({\boldsymbol{y}})_{q})+v_{1,p}\bigg{]}_{p=1}^{d_{2}}.$
    For any ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ and any index $p$, we have
    $\partial^{\boldsymbol{\nu}}G_{\theta}^{[1]}({\boldsymbol{y}})_{p}=\sum_{q=1}^{d_{1}}W_{1,p,q}\,\partial^{\boldsymbol{\nu}}\big{(}\sigma(G_{\theta}^{[0]}({\boldsymbol{y}})_{q})\big{)}.$ (86)
    For each index $q$, the recursive form of the multivariate Faà di Bruno formula gives (see [45, formula (3.1) and (3.5)] with the dimensionality $m$ there set to $1$)
    $\partial^{\boldsymbol{\nu}}\big{(}\sigma(G_{\theta}^{[0]}({\boldsymbol{y}})_{q})\big{)}$ $=\sum_{\lambda=1}^{|{\boldsymbol{\nu}}|}\sigma^{(\lambda)}(G_{\theta}^{[0]}({\boldsymbol{y}})_{q})\,\alpha_{{\boldsymbol{\nu}},\lambda}^{[0,q]}({\boldsymbol{y}}),$ (87)
    where the auxiliary functions $\alpha_{{\boldsymbol{\nu}},\lambda}^{[0,q]}({\boldsymbol{y}})$ are defined recursively by $\alpha_{{\boldsymbol{\nu}},0}^{[0,q]}({\boldsymbol{y}}):=\delta_{{\boldsymbol{\nu}},{\boldsymbol{0}}}$, $\alpha_{{\boldsymbol{\nu}},\lambda}^{[0,q]}({\boldsymbol{y}}):=0$ for $\lambda>|{\boldsymbol{\nu}}|$, and otherwise
    $\alpha_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}^{[0,q]}({\boldsymbol{y}}):=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\partial^{{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}G_{\theta}^{[0]}({\boldsymbol{y}})_{q}\,\alpha_{{\boldsymbol{\nu}}-{\boldsymbol{m}},\lambda-1}^{[0,q]}({\boldsymbol{y}}).$
    (Note that we have written the recursion differently from the non-periodic case, with ${\boldsymbol{m}}$ and ${\boldsymbol{\nu}}-{\boldsymbol{m}}$ swapped over, for convenience here.) Using (85), we obtain
    $|\alpha_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}^{[0,q]}({\boldsymbol{y}})|$ $\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}|\partial^{{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}G_{\theta}^{[0]}({\boldsymbol{y}})_{q}|\,|\alpha_{{\boldsymbol{\nu}}-{\boldsymbol{m}},\lambda-1}^{[0,q]}({\boldsymbol{y}})|$
    $\leq\beta_{j}\sum_{k=0}^{\nu_{j}}(2\pi)^{k+1}\,\binom{\nu_{j}}{k}\,|\alpha_{{\boldsymbol{\nu}}-k{\boldsymbol{e}}_{j},\lambda-1}^{[0,q]}({\boldsymbol{y}})|.$
    To solve this recurrence, we apply [19, Lemma A.4] (taking there $c=2\pi$ and $d=1$ which is the dimensionality of ${\boldsymbol{\lambda}}$ there) to obtain the bound
    $|\alpha_{{\boldsymbol{\nu}},\lambda}^{[0,q]}({\boldsymbol{y}})|$ $\leq(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}}}{{\scriptstyle{|{\boldsymbol{m}}|=\lambda}}}}{\boldsymbol{\beta}}^{\boldsymbol{m}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}).$
    (This bound is tight in the sense that it would be an equality if the recurrence were an equality.) Since this bound holds for all indices $q$, we apply this bound in (86) and (87) to yield for ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$,
    $|\partial^{\boldsymbol{\nu}}G_{\theta}^{[1]}({\boldsymbol{y}})_{p}|$ $\leq\sum_{q=1}^{d_{1}}|W_{1,p,q}|\,\sum_{\lambda=1}^{|{\boldsymbol{\nu}}|}A_{\lambda}\,(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}}}{{\scriptstyle{|{\boldsymbol{m}}|=\lambda}}}}{\boldsymbol{\beta}}^{\boldsymbol{m}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}})$
    $\leq(2\pi)^{|{\boldsymbol{\nu}}|}\,R_{1}\,\sum_{\lambda=1}^{|{\boldsymbol{\nu}}|}A_{\lambda}\,\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}}}{{\scriptstyle{|{\boldsymbol{m}}|=\lambda}}}}{\boldsymbol{\beta}}^{\boldsymbol{m}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}})$
    $=(2\pi)^{|{\boldsymbol{\nu}}|}\,R_{1}\,\sum_{{\boldsymbol{0}}\neq{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}{\boldsymbol{\beta}}^{\boldsymbol{m}}\,A_{|{\boldsymbol{m}}|}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}})=(2\pi)^{|{\boldsymbol{\nu}}|}\,R_{1}\,\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}{\boldsymbol{\beta}}^{\boldsymbol{m}}\,\Gamma_{|{\boldsymbol{m}}|}^{[1]}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}),$
    where we used the assumptions (8) and (9). The ${\boldsymbol{m}}={\boldsymbol{0}}$ term is 00 so it was added back into the sum. (This is because ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ so there exists a $\nu_{j}>0$ and ${\mathcal{S}}(\nu_{j},0)=0$.) We also used $A_{|{\boldsymbol{m}}|}=\Gamma_{|{\boldsymbol{m}}|}^{[1]}$ from (12). Hence (11) holds for $\ell=1$.
    Now consider $\ell\geq 2$ and suppose the result (11) holds for depth $\ell-1$. We can write
    $G_{\theta}^{[\ell]}({\boldsymbol{y}})=W_{\ell}\,\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}}))+{\boldsymbol{v}}_{\ell}=\bigg{[}\sum_{q=1}^{d_{\ell}}W_{\ell,p,q}\,\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q})+v_{\ell,p}\bigg{]}_{p=1}^{d_{\ell+1}}.$
    For any ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$ and any index $p$, we have
    $\partial^{\boldsymbol{\nu}}G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}$ $=\sum_{q=1}^{d_{\ell}}W_{\ell,p,q}\,\partial^{\boldsymbol{\nu}}\big{(}\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q})\big{)}.$ (88)
    For each index $q$, the recursive form of the multivariate Faà di Bruno formula gives (see [45, formula (3.1) and (3.5)] with the dimensionality $m$ there set to $1$)
    $\partial^{\boldsymbol{\nu}}\big{(}\sigma(G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q})\big{)}$ $=\sum_{\lambda=1}^{|{\boldsymbol{\nu}}|}\sigma^{(\lambda)}(G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q})\,\alpha_{{\boldsymbol{\nu}},\lambda}^{[\ell-1,q]}({\boldsymbol{y}}),$ (89)
    where the auxiliary functions $\alpha_{{\boldsymbol{\nu}},\lambda}^{[\ell-1,q]}({\boldsymbol{y}})$ are defined recursively by $\alpha_{{\boldsymbol{\nu}},0}^{[\ell-1,q]}({\boldsymbol{y}}):=\delta_{{\boldsymbol{\nu}},{\boldsymbol{0}}}$, $\alpha_{{\boldsymbol{\nu}},\lambda}^{[\ell-1,q]}({\boldsymbol{y}}):=0$ for $\lambda>|{\boldsymbol{\nu}}|$, and otherwise
    $\alpha_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}^{[\ell-1,q]}({\boldsymbol{y}}):=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\partial^{{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q}\,\alpha_{{\boldsymbol{m}},\lambda-1}^{[\ell-1,q]}({\boldsymbol{y}}).$
    Using the induction hypothesis (11) for depth $\ell-1$, we obtain
    $|\alpha_{{\boldsymbol{\nu}}+{\boldsymbol{e}}_{j},\lambda}^{[\ell-1,q]}({\boldsymbol{y}})|\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}|\partial^{{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}G_{\theta}^{[\ell-1]}({\boldsymbol{y}})_{q}|\,|\alpha_{{\boldsymbol{m}},\lambda-1}^{[\ell-1,q]}({\boldsymbol{y}})|$ (90)
    $\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\bigg{(}(2\pi)^{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|+1}\,R_{\ell-1}\,\sum_{{\boldsymbol{w}}\leq{\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j}}{\boldsymbol{\beta}}^{{\boldsymbol{w}}}\,\Gamma_{|{\boldsymbol{w}}|}^{[\ell-1]}\,{\mathcal{S}}({\boldsymbol{\nu}}-{\boldsymbol{m}}+{\boldsymbol{e}}_{j},{\boldsymbol{w}})\bigg{)}|\alpha_{{\boldsymbol{m}},\lambda-1}^{[\ell-1,q]}({\boldsymbol{y}})|.$
    Lemma 7 (with inequalities $\leq$) then yields for all ${\boldsymbol{\nu}}\neq{\boldsymbol{0}}$,
    $|\alpha_{{\boldsymbol{\nu}},\lambda}^{[\ell-1,q]}({\boldsymbol{y}})|\leq R_{\ell-1}^{\lambda}\,(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{u}}\leq{\boldsymbol{\nu}}}}}{{\scriptstyle{|{\boldsymbol{u}}|\geq\lambda}}}}{\boldsymbol{\beta}}^{{\boldsymbol{u}}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{u}}),$ (91)
    where the sequence ${\mathbb{B}}_{n,\lambda}$ is defined by (74) with $\Gamma_{n}=\Gamma_{n}^{[\ell-1]}$, and this sequence ${\mathbb{B}}_{n,\lambda}$ is precisely ${\mathbb{B}}_{n,\lambda}^{[\ell-1]}$ defined in (13). Since the same upper bound (91) holds for all indices $q$, we conclude from (88) and (89) that
    $|\partial^{\boldsymbol{\nu}}G_{\theta}^{[\ell]}({\boldsymbol{y}})_{p}|$ $\leq\sum_{q=1}^{d_{\ell}}|W_{\ell,p,q}|\sum_{\lambda=1}^{|{\boldsymbol{\nu}}|}A_{\lambda}\,R_{\ell-1}^{\lambda}\,(2\pi)^{|{\boldsymbol{\nu}}|}\,\sum_{\stackrel{{\scriptstyle\scriptstyle{{\boldsymbol{u}}\leq{\boldsymbol{\nu}}}}}{{\scriptstyle{|{\boldsymbol{u}}|\geq\lambda}}}}{\boldsymbol{\beta}}^{{\boldsymbol{u}}}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda}^{[\ell-1]}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{u}})$
    $\leq(2\pi)^{|{\boldsymbol{\nu}}|}\,R_{\ell}\,\sum_{{\boldsymbol{u}}\leq{\boldsymbol{\nu}}}{\boldsymbol{\beta}}^{{\boldsymbol{u}}}\bigg{(}\underbrace{\sum_{\lambda=1}^{|{\boldsymbol{u}}|}A_{\lambda}\,R_{\ell-1}^{\lambda}\,{\mathbb{B}}_{|{\boldsymbol{u}}|,\lambda}^{[\ell-1]}}_{=\,\Gamma_{|{\boldsymbol{u}}|}^{[\ell]}}\bigg{)}{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{u}}),$
[…]
6.3 Proofs for popular activation functions
The logistic sigmoid function
$\sigma(x)=\frac{1}{1+e^{-x}}$ (92)
is monotone on ${\mathbb{R}}$ and has range $(0,1)$. From [35, formula (15)] we have for $n\geq 1$
$\sigma^{(n)}(x)=\sum_{k=1}^{n}(-1)^{k-1}E(n,k-1)\,[\sigma(x)]^{k}\,[1-\sigma(x)]^{n+1-k},$ (93)
where $E(n,k):=\sum_{i=0}^{k}(-1)^{i}\binom{n+1}{i}(k+1-i)^{n}$ is the Eulerian number, i.e., the number of permutations of the integers $1$ to $n$ in which exactly $k$ elements are greater than the previous element. Thus
$|\sigma^{(n)}(x)|\leq\sum_{k=1}^{n}E(n,k-1)=\sum_{k=0}^{n-1}E(n,k)=n!\,,$
where conveniently the sum of Eulerian numbers is equal to $n!$
We can actually obtain a tighter upper bound. Indeed, the function $f_{k}(y):=y^{k}(1-y)^{n+1-k}$ for $1\leq k\leq n$ vanishes when $y=0$ and $y=1$; it is positive on the interval $(0,1)$ with a unique maximum at $y=\frac{k}{n+1}$. Furthermore, $f_{k}(\frac{k}{n+1})$ is largest when $k=1$ or $n$, and this value is $f_{1}(\frac{1}{n+1})=\frac{n^{n}}{(n+1)^{n+1}}$. Thus for all $n\geq 1$ we have
$|\sigma^{(n)}(x)|\leq f_{1}\Big{(}\frac{1}{n+1}\Big{)}\sum_{k=1}^{n}E(n,k-1)=\frac{n^{n}}{(n+1)^{n+1}}\sum_{k=0}^{n-1}E(n,k)$ $=\frac{n^{n}}{(n+1)^{n+1}}\,n!\,.$ (94)
The hyperbolic tangent activation function
is related to the logistic sigmoid (92) by
$\tanh(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}=2\,\sigma(2x)-1,$
and therefore for $n\geq 1$ we have
$\frac{{\mathrm{d}}^{n}}{{\mathrm{d}}x^{n}}\tanh(x)=2^{n+1}\,\sigma^{(n)}(2x),$
which combined with (94) gives
$\Big{|}\frac{{\mathrm{d}}^{n}}{{\mathrm{d}}x^{n}}\tanh(x)\Big{|}\leq 2^{n+1}\,\frac{1}{n}\Big{(}\frac{n}{n+1}\Big{)}^{n+1}n!<2^{n}\,n!\qquad\mbox{for all $n\geq 1$.}$
The swish activation function
is related to the logistic sigmoid function (92) by
${\rm swish}(x)=\frac{x}{1+e^{-x}}=x\,\sigma(x),$
and therefore
$\frac{{\mathrm{d}}^{n}}{{\mathrm{d}}x^{n}}{\rm swish}(x)=x\,\sigma^{(n)}(x)+n\,\sigma^{(n-1)}(x).$
Using (93) we obtain
$x\,\sigma^{(n)}(x)$ $=\sum_{k=1}^{n}(-1)^{k-1}E(n,k-1)\,x\,[\sigma(x)]^{k}\,[1-\sigma(x)]^{n+1-k}$
$=\sum_{k=1}^{n}(-1)^{k-1}E(n,k-1)\,g(x)\,[\sigma(x)]^{k-1}\,[1-\sigma(x)]^{n-k},$
where (noting that $k-1\geq 0$ and $n-k\geq 0$) we defined
$g(x):=x\,\sigma(x)\,(1-\sigma(x))=\frac{x\,e^{-x}}{(1+e^{-x})^{2}}=\frac{x}{(e^{x/2}+e^{-x/2})^{2}}.$
Clearly $g$ is an odd function and $g^{\prime}(x)=e^{-x}(1-x+e^{-x}+x\,e^{-x})/(1+e^{-x})^{3}$ vanishes exactly when $x\approx\pm 1.543$, leading to two global extrema, and therefore $|g(x)|\leq g_{\max}\approx 0.224$. Thus
$|x\,\sigma^{(n)}(x)|\leq g_{\max}\sum_{k=1}^{n}E(n,k-1)=g_{\max}\sum_{k=0}^{n-1}E(n,k)=g_{\max}\,n!\,.$ (95)
Hence we can use (95) and (94) to obtain for $n\geq 2$,
$\Big{|}\frac{{\mathrm{d}}^{n}}{{\mathrm{d}}x^{n}}{\rm swish}(x)\Big{|}$ $\leq|x\,\sigma^{(n)}(x)|+n\,|\sigma^{(n-1)}(x)|$
$\leq g_{\max}\,n!+n\,\frac{(n-1)^{n-1}}{n^{n}}(n-1)!$
$=g_{\max}\,n!+\frac{1}{n}\Big{(}\frac{n-1}{n}\Big{)}^{n-1}\,n!<\Big{(}\frac{1}{4}+\frac{1}{n}\Big{)}n!<n!\,.$
For $n=1$ we have
$\frac{{\mathrm{d}}}{{\mathrm{d}}x}{\rm swish}(x)=x\,\sigma^{\prime}(x)+\sigma(x)=x\,\frac{e^{-x}}{(1+e^{-x})^{2}}+\frac{1}{1+e^{-x}}=\frac{1+e^{-x}+x\,e^{-x}}{(1+e^{-x})^{2}},$
and
$\frac{{\mathrm{d}}^{2}}{{\mathrm{d}}x^{2}}{\rm swish}(x)$ $=\frac{-x\,e^{-x}}{(1+e^{-x})^{2}}+\frac{2\,e^{-x}(1+e^{-x}+x\,e^{-x})}{(1+e^{-x})^{3}}$
$=\frac{e^{-x}[-x(1+e^{-x})+2(1+e^{-x}+x\,e^{-x})]}{(1+e^{-x})^{3}}=\frac{e^{-x}[-x+2+2e^{-x}+x\,e^{-x}]}{(1+e^{-x})^{3}};$
the latter vanishes at $x\approx\pm 2.3994$, giving local maximum/mininum values $1.0998$ and $-0.0998$ for the first derivative. Thus $|\frac{{\mathrm{d}}}{{\mathrm{d}}x}{\rm swish}(x)|\leq 1.1$. Combining this with the upper bound $n!$ for $n\geq 2$, we conclude that
$\Big{|}\frac{{\mathrm{d}}^{n}}{{\mathrm{d}}x^{n}}{\rm swish}(x)\Big{|}\leq 1.1\,n!\qquad\mbox{for all $n\geq 1$.}$
In summary
we have shown that the bound $A_{n}$ in (9) can be taken as
$A_{n}=\begin{cases}n!&\text{for sigmoid},\\ 2^{n}\,n!&\text{for tanh},\\ 1.1\,n!&\text{for swish},\end{cases}$ (96)
leading us to consider the general form $A_{n}=\xi\,\tau^{n}\,n!$
Lemma 8
Let $\xi>0$ and $\tau>0$. If $A_{n}=\xi\,\tau^{n}\,n!$ for $n\geq 1$ and $R_{\ell}>0$ for $\ell\geq 1$, then the sequences defined recursively in (12)–(13) are given explicitly by
$\Gamma_{n}^{[\ell]}={\mathbb{B}}_{n,1}^{[\ell]}$ $=P_{\ell-1}\,\bigg{(}\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{n-1}\,\xi\,\tau^{n}\,n!\qquad\mbox{for $\ell\geq 1$ and $n\geq 1$},$ (97)
${\mathbb{B}}_{n,\lambda}^{[\ell]}$ $=P_{\ell-1}^{\lambda}\,\bigg{(}\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{n-\lambda}\,\xi^{\lambda}\,\tau^{n}\,\frac{n!}{\lambda!}\binom{n-1}{\lambda-1}\qquad\mbox{for $\ell\geq 1$ and $n\geq\lambda\geq 1$},$ (98)
where
$P_{0}:=1,\qquad P_{k}:=\prod_{t=1}^{k}(\xi\,\tau R_{t})\qquad\mbox{for $k\geq 1$}.$ (99)
Proof.
For $\ell=1$, we have from (12)–(13) that $\Gamma_{n}^{[1]}={\mathbb{B}}_{n,1}^{[1]}=A_{n}=\xi\,\tau^{n}\,n!$ for all $n\geq 1$, which agrees with (97) since $P_{0}=1$ from (99). Thus the formula (97) holds for $\ell=1$ and all $n\geq 1$. And trivially, the formula (98) holds for $\ell=1$ and $\lambda=1$ and all $n\geq\lambda=1$.
For $\ell=1$ and some $\lambda\geq 2$, assume that the formula (98) for ${\mathbb{B}}_{n,\lambda^{\prime}}^{[1]}$ holds for all $1\leq\lambda^{\prime}<\lambda$ and all $n\geq\lambda^{\prime}$. Then from the recursion (13) we have
${\mathbb{B}}_{n,\lambda}^{[1]}=\sum_{i=\lambda-1}^{n-1}\binom{n-1}{i}\,\Gamma_{n-i}^{[1]}\,{\mathbb{B}}_{i,\lambda-1}^{[1]}$ $=\sum_{i=\lambda-1}^{n-1}\binom{n-1}{i}\,\xi\,\tau^{n-i}\,(n-i)!\,\xi^{\lambda-1}\,\tau^{i}\,\frac{i!}{(\lambda-1)!}\binom{i-1}{\lambda-2}$
$=\xi^{\lambda}\,\tau^{n}\,\frac{(n-1)!}{(\lambda-1)!}\,\sum_{i=\lambda-1}^{n-1}(n-i)\binom{i-1}{\lambda-2},$
with
$\sum_{i=\lambda-1}^{n-1}(n-i)\,\binom{i-1}{\lambda-2}$ $=n\sum_{i=\lambda-1}^{n-1}\binom{i-1}{\lambda-2}-(\lambda-1)\sum_{i=\lambda-1}^{n-1}\binom{i}{\lambda-1}$
$=n\sum_{i=\lambda-2}^{n-2}\binom{i}{\lambda-2}-(\lambda-1)\sum_{i=\lambda-1}^{n-1}\binom{i}{\lambda-1}$
$=n\binom{n-1}{\lambda-1}-(\lambda-1)\binom{n}{\lambda}=\lambda\binom{n}{\lambda}-(\lambda-1)\binom{n}{\lambda}=\binom{n}{\lambda},$ (100)
where we used the “hockey-stick identity” $\sum_{m=k}^{n}\binom{m}{k}=\binom{n+1}{k+1}$. Thus we have
${\mathbb{B}}_{n,\lambda}^{[1]}=\xi^{\lambda}\,\tau^{n}\,\frac{(n-1)!}{(\lambda-1)!}\,\binom{n}{\lambda}=\xi^{\lambda}\,\tau^{n}\,\frac{n!}{\lambda!}\,\binom{n-1}{\lambda-1}.$ (101)
Hence by induction on $\lambda$, the formula (98) holds for $\ell=1$ and all $n\geq\lambda\geq 1$.
Consider now a fixed level $\ell\geq 2$ and assume that the formulas (97)–(98) hold in level $\ell-1$. Then for all $n\geq 1$ we have from (12) that
$\Gamma_{n}^{[\ell]}={\mathbb{B}}_{n,1}^{[\ell]}=\sum_{\lambda=1}^{n}A_{\lambda}\,R_{\ell-1}^{\lambda}\,{\mathbb{B}}_{n,\lambda}^{[\ell-1]}$ $=\sum_{\lambda=1}^{n}\xi\,\tau^{\lambda}\,\lambda!\,R_{\ell-1}^{\lambda}\,P_{\ell-2}^{\lambda}\,\bigg{(}\sum_{k=0}^{\ell-2}P_{k}\bigg{)}^{n-\lambda}\,\xi^{\lambda}\,\tau^{n}\,\frac{n!}{\lambda!}\,\binom{n-1}{\lambda-1}$
$=\xi\,\tau^{n}\,n!\,\sum_{\lambda=1}^{n}P_{\ell-1}^{\lambda}\,\bigg{(}\sum_{k=0}^{\ell-2}P_{k}\bigg{)}^{n-\lambda}\,\binom{n-1}{\lambda-1}$
$=\xi\,\tau^{n}\,n!\,P_{\ell-1}\sum_{\lambda^{\prime}=0}^{n-1}P_{\ell-1}^{\lambda^{\prime}}\,\bigg{(}\sum_{k=0}^{\ell-2}P_{k}\bigg{)}^{n-1-\lambda^{\prime}}\,\binom{n-1}{\lambda^{\prime}}$
$=\xi\,\tau^{n}\,n!\,P_{\ell-1}\,\bigg{(}P_{\ell-1}+\sum_{k=0}^{\ell-2}P_{k}\bigg{)}^{n-1}=P_{\ell-1}\bigg{(}\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{n-1}\,\xi\,\tau^{n}\,n!\,,$
where in the fourth equality we used $(\xi\,\tau\,R_{\ell-1})\,P_{\ell-2}=P_{\ell-1}$ from (99), and in the sixth equality we used the binomial theorem. Thus the formula (98) holds for this level $\ell\geq 2$ and all $n\geq 1$. Also trivially, the formula (98) holds for this level $\ell\geq 2$ and $\lambda=1$, and all $n\geq\lambda$, because (98) coincides with (97) when $\lambda=1$.
For the same level $\ell\geq 2$ and some $\lambda\geq 2$, assume that the formula (98) for ${\mathbb{B}}_{n,\lambda^{\prime}}^{[\ell]}$ holds for all $1\leq\lambda^{\prime}<\lambda$ and all $n\geq\lambda^{\prime}$. Then from the recursion (13) we have
${\mathbb{B}}_{n,\lambda}^{[\ell]}=\sum_{i=\lambda-1}^{n-1}\binom{n-1}{i}\,\Gamma_{n-i}^{[\ell]}\,{\mathbb{B}}_{i,\lambda-1}^{[\ell]}$ $=\sum_{i=\lambda-1}^{n-1}\binom{n-1}{i}\,P_{\ell-1}\,\bigg{(}\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{n-i-1}\,\xi\,\tau^{n-i}\,(n-i)!$
$\qquad\qquad\qquad\cdot P_{\ell-1}^{\lambda-1}\,\bigg{(}\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{i-(\lambda-1)}\,\xi^{\lambda-1}\,\tau^{i}\,\frac{i!}{(\lambda-1)!}\binom{i-1}{\lambda-2}$
$=P_{\ell-1}^{\lambda}\,\bigg{(}\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{n-\lambda}\,\xi^{\lambda}\,\tau^{n}\,\frac{(n-1)!}{(\lambda-1)!}\sum_{i=\lambda-1}^{n-1}(n-i)\,\binom{i-1}{\lambda-2}$
$=P_{\ell-1}^{\lambda}\,\bigg{(}\sum_{k=0}^{\ell-1}P_{k}\bigg{)}^{n-\lambda}\,\xi^{\lambda}\,\tau^{n}\,\frac{n!}{\lambda!}\,\binom{n-1}{\lambda-1},$
where the last step follows from (6.3) and the second equality in (101). This is exactly (98), thus by induction on $\lambda$, the formula (98) holds for this level $\ell\geq 2$ and all $n\geq\lambda\geq 1$.
Hence by induction on the level $\ell$, we conclude that the formulas (97)–(98) hold for all levels $\ell\geq 1$ and all $n\geq\lambda\geq 1$. This completes the proof. ∎
6.4 Proofs for bounds on norms
  • Proof of Theorem 4.
    Consider first the non-periodic case. With the restrictions (36)–(39), both the target function $G({\boldsymbol{y}})$ and the DNN $G^{[L]}_{\theta}({\boldsymbol{y}})$ satisfy the regularity bound (40), and thus their difference satisfies
    $\big{|}\partial^{{\boldsymbol{\nu}}}\big{(}G({\boldsymbol{y}})_{p}-G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{)}\big{|}\leq\big{|}\partial^{{\boldsymbol{\nu}}}G({\boldsymbol{y}})_{p}\big{|}+\big{|}\partial^{{\boldsymbol{\nu}}}G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{|}\leq 2\,C\,|{\boldsymbol{\nu}}|!\,{\boldsymbol{b}}^{\boldsymbol{\nu}}.$
    For the square of the difference, we use the Leibniz product rule to obtain
    $\big{|}\partial^{{\boldsymbol{\nu}}}\big{(}G({\boldsymbol{y}})_{p}-G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{)}^{2}\big{|}$ $\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\big{|}\partial^{{\boldsymbol{m}}}\big{(}G({\boldsymbol{y}})_{p}-G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{)}\big{|}\,\big{|}\partial^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}\big{(}G({\boldsymbol{y}})_{p}-G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{)}\big{|}$
    $\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\big{(}2\,C\,|{\boldsymbol{m}}|!\,{\boldsymbol{b}}^{\boldsymbol{m}}\big{)}\big{(}2\,C\,|{\boldsymbol{\nu}}-{\boldsymbol{m}}|!\,{\boldsymbol{b}}^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}\big{)}$
    $=4\,C^{2}\,(|{\boldsymbol{\nu}}|+1)!\,{\boldsymbol{b}}^{\boldsymbol{\nu}}\,,$
    where we used the identity (see e.g., [30, formula (9.4)])
    $\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\,|{\boldsymbol{m}}|!\,|{\boldsymbol{\nu}}-{\boldsymbol{m}}|!=(|{\boldsymbol{\nu}}|+1)!\,.$ (102)
    This yields the norm bound (44) as required.
    Analogously, for the periodic case both the target function $G({\boldsymbol{y}})$ and the DNN $G^{[L]}_{\theta}({\boldsymbol{y}})$ satisfy the regularity bound (45). Thus we have
    $\big{|}\partial^{{\boldsymbol{\nu}}}\big{(}G({\boldsymbol{y}})_{p}-G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{)}\big{|}\leq\big{|}\partial^{{\boldsymbol{\nu}}}G({\boldsymbol{y}})_{p}\big{|}+\big{|}\partial^{{\boldsymbol{\nu}}}G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{|}\leq 2\,C\,(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}|{\boldsymbol{m}}|!\,{\boldsymbol{b}}^{\boldsymbol{m}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}),$
    and
    $\big{|}\partial^{{\boldsymbol{\nu}}}\big{(}G({\boldsymbol{y}})_{p}-G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{)}^{2}\big{|}\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\big{|}\partial^{{\boldsymbol{m}}}\big{(}G({\boldsymbol{y}})_{p}-G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{)}\big{|}\,\big{|}\partial^{{\boldsymbol{\nu}}-{\boldsymbol{m}}}\big{(}G({\boldsymbol{y}})_{p}-G^{[L]}_{\theta}({\boldsymbol{y}})_{p}\big{)}\big{|}$
    $\leq\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\bigg{(}2\,C\,(2\pi)^{|{\boldsymbol{m}}|}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{m}}}|{\boldsymbol{w}}|!\,{\boldsymbol{b}}^{\boldsymbol{w}}\,{\mathcal{S}}({\boldsymbol{m}},{\boldsymbol{w}})\!\bigg{)}\bigg{(}2\,C\,(2\pi)^{|{\boldsymbol{\nu}}-{\boldsymbol{m}}|}\sum_{{\boldsymbol{u}}\leq{\boldsymbol{\nu}}-{\boldsymbol{m}}}|{\boldsymbol{u}}|!\,{\boldsymbol{b}}^{\boldsymbol{u}}\,{\mathcal{S}}({\boldsymbol{\nu}}-{\boldsymbol{m}},{\boldsymbol{u}})\!\bigg{)}$
    $=4\,C^{2}\,(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\bigg{(}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{m}}}\binom{{\boldsymbol{m}}}{{\boldsymbol{w}}}\,|{\boldsymbol{w}}|!\,|{\boldsymbol{m}}-{\boldsymbol{w}}|!\bigg{)}{\boldsymbol{b}}^{{\boldsymbol{m}}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}})$
    $=4\,C^{2}\,(2\pi)^{|{\boldsymbol{\nu}}|}\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}(|{\boldsymbol{m}}|+1)!\,{\boldsymbol{b}}^{{\boldsymbol{m}}}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}}),$
    where in the penultimate step we used the identity [19, Lemma A.3]: for any ${\boldsymbol{\nu}}\in\mathcal{I}$ and any arbitrary sequences $({\mathbb{A}}_{\boldsymbol{\nu}})_{{\boldsymbol{\nu}}\in\mathcal{I}}$ and $({\mathbb{B}}_{\boldsymbol{\nu}})_{{\boldsymbol{\nu}}\in\mathcal{I}}$ of real numbers, we have
    $\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\binom{{\boldsymbol{\nu}}}{{\boldsymbol{m}}}\bigg{(}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{m}}}{\mathbb{A}}_{\boldsymbol{w}}\,{\mathcal{S}}({\boldsymbol{m}},{\boldsymbol{w}})\bigg{)}\,\bigg{(}\sum_{{\boldsymbol{u}}\leq{\boldsymbol{\nu}}-{\boldsymbol{m}}}{\mathbb{B}}_{\boldsymbol{\nu}}\,{\mathcal{S}}({\boldsymbol{\nu}}-{\boldsymbol{m}},{\boldsymbol{u}})\bigg{)}$
    $=\sum_{{\boldsymbol{m}}\leq{\boldsymbol{\nu}}}\bigg{(}\sum_{{\boldsymbol{w}}\leq{\boldsymbol{m}}}\binom{{\boldsymbol{m}}}{{\boldsymbol{w}}}{\mathbb{A}}_{\boldsymbol{w}}\,{\mathbb{B}}_{{\boldsymbol{m}}-{\boldsymbol{w}}}\bigg{)}\,{\mathcal{S}}({\boldsymbol{\nu}},{\boldsymbol{m}});$
[…]
7 Conclusion
We obtain new explicit regularity bounds (Theorem 1) for both the standard non-periodic DNN (2) and our newly proposed periodic DNN (4). The bounds depend on the neural network parameters as well as the choice of smooth activation function. We see that three important factors need to be under control: the columns of the matrix $W_{0}$ need to decay (cf. the sequence $\beta_{j}$ (7)), the matrices $W_{1},\ldots,W_{L-1}$ need to be bounded (cf. the sequence $R_{\ell}$ (8)), and the derivatives of the activation function need to be bounded (cf. the sequence $A_{n}$ (9)). We also state regularity bounds in the form that covers three commonly used activation functions: sigmoid, tanh, and swish (Theorem 2).
We provide bounds of the mixed-derivative-based norms of the DNNs (Theorem 3). By imposing restrictions on the network parameters to match the regularity features of the target functions (Theorem 4), we prove that DNNs with $N$ tailor-constructed lattice training points can achieve the generalization error or $L_{2}$ approximation error bound ${\tt tol}+{\mathcal{O}}(N^{-r/2})$, where ${\tt tol}\in(0,1)$ is the tolerance achieved by the training error in practice, and $r$ is related in different ways depending on the settings (Theorem 5) to the summability exponent $p^{*}$ of a sequence that characterises the decay of the input variables in the target functions, and with the implied constant independent of the dimensionality of the input data.
We restrict the network parameters during training by adding tailored regularization terms. For an algebraic equation mimicking the parametric PDE problems, our simple numerical experiments with the periodic DNNs and the sigmoid activation function for a periodic algebraic target function show that the DNNs trained with tailored regularization perform significantly better than the standard $\ell_{2}$ regularization. More detailed experiments, including with other activation functions, are considered in our survey article [28].
Acknowledgements
We acknowledge the financial support from the Australian Research Council Discovery Project DP240100769.